### v1

# Safe In-Context Reinforcement Learning

Abstract
In-context reinforcement learning (ICRL) is an emerging RL paradigm where the agent, after some pretraining procedure, is able to adapt to out-of-distribution test tasks without any parameter updates. The agent achieves this by continually expanding the input (i.e., the context) to its policy neural networks. For example, the input could be all the history experience that the agent has access to until the current time step. The agent’s performance improves as the input grows, without any parameter updates. In this work, we propose the first method that promotes the safety of ICRL’s adaptation process in the framework of constrained Markov Decision Processes. In other words, during the parameter-update-free adaptation process, the agent not only maximizes the reward but also minimizes an additional cost function. We also demonstrate that our agent actively reacts to the threshold (i.e., budget) of the cost tolerance. With a higher cost budget, the agent behaves more aggressively, and with a lower cost budget, the agent behaves more conservatively.
1 Introduction
Reinforcement learning (RL, Sutton & Barto (2018)) is a framework for solving sequential decision making problems via trial and error. Modern RL agents are usually parameterized by deep neural networks, and the trial-and-error process is typically achieved by updating the parameters of the neural networks (Mnih et al., 2015). Recently, in-context reinforcement learning (ICRL, Moeini et al. (2025)) emerges as an RL paradigm that achieves this trial-and-error process without any neural network parameter updates (Duan et al., 2016; Mishra et al., 2018; Xu et al., 2022; Kirsch et al., 2023; Lin et al., 2023; Lu et al., 2023; Raparthy et al., 2023; Sinii et al., 2023; Grigsby et al., 2024a; b; Cook et al., 2024; Dai et al., 2024; Dong et al., 2024; Elawady et al., 2024; Huang et al., 2024; Krishnamurthy et al., 2024; Laskin et al., 2023; Lee et al., 2023; Shi et al., 2024; Wang et al., 2025a; Zisman et al., 2023; 2024; Song et al., 2025; Wang et al., 2025b)
Specifically, the pretrained agent neural network takes as input not only the observation at the current time step but also an additional context. For example, the simplest choice of the context can be all the history of experiences that the agent has access to up to the current time step (Laskin et al., 2023). It is then observed that the agent’s performance improves as the length of the context grows, even if the neural network parameters are kept fixed and even if the agent is evaluated on a task that it has never encountered during the pretraining process (Laskin et al., 2023). This phenomenon is called ICRL. A widely accepted hypothesis for this performance improvement is that the forward pass of the pretrained agent network implicitly implements some RL algorithm during inference time to process the information in the context (Laskin et al., 2023; Lin et al., 2023; Wang et al., 2025a; b). As the context length grows, the amount of information grows, and the implemented RL algorithm in the forward pass is then able to output better actions.
ICRL has recently received increasing attention given its cheap yet strong adaptation capability. The adaptation is cheap in that it does not require parameter tuning. So to support such adaptation, the infrastructure only needs to support the inference of the neural network. For example, by using ICRL in LLMs’ inference time, the LLMs are able to significantly improve on challenging tasks such as ScienceWorld (Wang et al., 2022), HMMT (Harvard–MIT Mathematics Tournament, ), and AIME (Mathematical Association of America, ) without any fine-tuning (Song et al., 2025). The adaptation is strong in that ICRL allows adaptation to not only unseen but also out-of-distribution tasks. For example, Laskin et al. (2023) show that after pretraining on some bandit tasks, their ICRL agent is able to adapt quickly to test bandit tasks that have opposite optimal arms to the pretraining bandit tasks. Kirsch et al. (2023) show that after pretraining on Ant robot manipulation tasks, their ICRL agent is able to adapt to pole balancing tasks.
Despite the empirical success and the promising potential, the safety of ICRL has long been ignored. Here we discuss safety in the framework of constrained Markov Decision Processes (Altman, 2021), where each action the agent takes not only incurs a reward but also incurs a cost. Then when an agent adapts to a new task, it needs to not only maximize the rewards but also minimize the cost, all without updating the parameters.
We argue that safe ICRL is a necessary condition to enable many possible real-world applications of ICRL, such as embodied AI. Although such safe decision-making problems have been well studied in the classical RL setup (Garcıa & Fernández, 2015), to our knowledge, no prior work has studied the safety of ICRL, which is the gap that this paper shall close. We envision that a safe ICRL agent should have the following two desiderata, both of which should be achieved without any parameter updates during the adaptation to new tasks.
  • When evaluated in out-of-distribution tasks, the agent should be able to maximize the rewards while keeping the cost below a user-specified threshold.
  • The agent should actively react to the cost threshold. For example, a smaller cost threshold should induce a more conservative behavior, while a larger cost threshold should induce a more aggressive behavior.
The key contribution of this work is to design the first safe ICRL agent that simultaneously fulfills the two desiderata. Specifically, we make the following two contributions.
  • We study both off-line (cf. (Laskin et al., 2023)) and on-line pretraining approaches (cf. (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b)) for safe ICRL. We identify key limitations of the offline approach and propose a novel online pretraining approach that fulfills the two desiderata.
  • We empirically validate our proposed approach in challenging safe decision making problems. Particularly, our test task is not only unseen during the pretraining stage but also out of the pretraining task distribution. As a result, to succeed in our benchmarks, the agent has to demonstrate strong out-of-distribution generalization. Particularly, the agent has to explore efficiently in out-of-distribution test tasks, all without parameter updates.
2 Background
We consider a finite horizon Markov Decision Process (MDP, Bellman (1957); Puterman (2014)) with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1)$, and a horizon length $T$. At time step 0, an initial state $S_{0}$ is sampled from $p_{0}$. At time step $t$, an agent at a state $S_{t}$ selects an action $A_{t}$ according to its policy $\pi:\mathcal{A}\times\mathcal{S}\to[0,1]$, i.e., $A_{t}\sim\pi(\cdot|S_{t})$. The agent then proceeeds to a successor state $S_{t+1}\sim p(\cdot|S_{t},A_{t})$ and obtains a reward $R_{t+1}\doteq r(S_{t},A_{t})$. This process continues until the time T1T-1, where the agent executes the action $A_{T-1}$ and receives the last reward $R_{T}$. We use $\tau\doteq(S_{0},A_{0},R_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T})$ to denote an episode. We use $k$ to index episodes and $t$ to index time steps within an episode. When necessary, we write $S_{t}^{k}$ to denote the state at time step $t$ within the $k$-th episode. $A_{t}^{k}$ and $R_{t}^{k}$ are similarly defined. For any episode $\tau$, we use $G(\tau)\doteq\sum_{t=1}^{T}R_{t}$ to denote the return associated with this episode. The performance of a policy $\pi$ is then measured by the expected total rewards $J(\pi)\doteq\mathbb{E}_{\tau\sim\pi}[G(\tau)]$. One fundamental task in RL, called control, is to adapt the policy $\pi$ to maximize $J(\pi)$.
Modern RL methods solve the control problem via parameterizing the policy $\pi$ with neural networks (Mnih et al., 2015). We denote such parameterized policy as $\pi_{\theta}$, where $\theta$ denotes the parameters of the neural network. Typically, $\theta$ is updated at every time step, e.g., $\theta_{t+1}$ is obtained by updating $\theta_{t}$ with the newly obtained information $(S_{t},A_{t},S_{t+1},R_{t+1})$. The performance of the policy $\pi_{\theta_{t}}$ improves along with such parameter updates. In other words, the reinforcement learning process occurs with the parameter updates. ICRL is an emerging paradigm that allows the reinforcement learning process to occur without any parameter updates. In ICRL, the policy $\pi$ takes an additional context as input. We denote the context at time $t$ within the $k$-th episode as $H_{t}^{k}$ and express the policy as $\pi(\cdot|S_{t}^{k},H_{t}^{k})$ to emphasize such dependency. The construction of the context is an active research area and we refer the reader to Moeini et al. (2025) for a detailed survey. The simplest example is to use all the history as context, e.g., $H_{t}^{k}\doteq(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},S_{1}^{k},\dots S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k})$. After some pretraining procedure on multiple MDPs to be detailed soon, we can obtain a parameter $\theta_{*}$. It is observed that the performance of $\pi_{\theta_{*}}(\cdot|S_{t}^{k},H_{t}^{k})$ improves as the context grows with $\theta_{*}$ kept fixed. In other words, the reinforcement learning process now occurs without any parameter updates. This cannot be attributed to the hypothesis that $\theta_{*}$ memorizes a good policy, as such performance improvement is also observed even when $\pi_{\theta_{*}}$ is evaluated in out-of-distribution MDPs that are very different from the pretraining MDPs (Laskin et al., 2023; Kirsch et al., 2023). A widely accepted hypothesis is that the neural network $\theta_{*}$ implicitly implements some RL algorithm in its forward pass, such that it can process the context with the RL algorithm in the inference time (Moeini et al., 2025). As the context grows, the inference-time RL algorithm in the forward pass gains access to more information. So the performance of the policy improves. This hypothesis is also theoretically backed by Lin et al. (2023); Wang et al. (2025a; b).
We now elaborate on the pretraining methods for ICRL. According to the taxonomy in Moeini et al. (2025), the pretraining can be divided into supervsied pretraining (Laskin et al., 2023; Lin et al., 2023) and reinforcement pretraining (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b; Wang et al., 2025a; b). Supervised pretraining is essentially behavior cloning but the goal is to imitate an algorithm instead of a policy. Supervised pretraining is usually done in an offline manner. Specifically, by running an existing RL algorithm on an MDP, we can collect a sequence of episodes, denoted as $\Xi\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$. We call $\Xi$ a trajectory. Suppose the RL algorithm behaves well, we would expect that the episode return $G(\tau_{k})$ increases as $k$ increases. So the trajectory $\Xi$ demonstrates the learning process of the RL algorithm in the MDP. In supervised pretraining, we collect multiple trajectories by running multiple existing RL algorithms on multiple MDPs, yielding a dataset $\quantity{\Xi_{i}}$. The policy $\pi_{\theta}$ is then trained with an imitation learning loss. In other words, for a state $S_{t}^{k}$ from a trajectory $\Xi_{i}$ in the dataset, the loss for updating $\theta$ is
$-\log\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k}).$ (1)
By asking the neural network to imitate the behavior of RL algorithms demonstrated in the dataset, we expect the neural network to be able to distill some RL algorithm into its forward pass. This is known as algorithm distillation (Laskin et al., 2023). Instead of distilling existing RL algorithms, reinforcement pretraining asks the network to discover its own RL algorithm. Reinforcement pretraining is typically done in an online manner. At time step $t$ within the $k$-th episode, the loss is
$Loss_{RL}(\pi_{\theta}(\cdot|S_{t}^{k},H_{t}^{k})),$ (2)
where $Loss_{RL}$ can be any standard RL loss for standard online RL algorithms, e.g., Grigsby et al. (2024a) use a variant of DDPG (Lillicrap et al., 2016), Elawady et al. (2024) use a variant of PPO (Schulman et al., 2017), and Cook et al. (2024) use Muesli (Hessel et al., 2021). Since the context is typically a long sequence, Transformers (Vaswani et al., 2017) or state space models (Gu & Dao, 2024) are usually used to parameterize the policy (Laskin et al., 2023; Lu et al., 2023). The long sequence context is one of the main driving forces for the remarkable generalization capability of ICRL.
3 Safe In-Context Reinforcement Learning
While ICRL has demonstrated remarkable generalization in out-of-distribution tasks, the safety of such generalization has been largely overlooked. This paper is the first to study the safety of ICRL. Particularly, we use the constrained MDP (CMDP) framework (Altman, 2021). In addition to the reward function $r$, we also have a cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. At a state $S_{t}$, the agent executes an action $A_{t}$ and receives both a reward $R_{t+1}$ and a cost $C_{t+1}\doteq c(S_{t},A_{t})$. An episode is then
$\tau=(S_{0},A_{0},R_{1},C_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T},C_{T}).$ (3)
We use $G_{c}(\tau)\doteq\sum_{t=1}^{T}C_{t}$ to denote the toal cost in the episode $\tau$. After some pretraining procedure, we obtain a policy $\pi_{\theta_{*}}$. We then execute this policy in a test MDP for multiple episodes $\tau_{1},\dots,\tau_{K}$. With ICRL, we can expect that the episode return $G(\tau_{k})$ increaes as $k$ increaes. But for safe ICRL, we would like to additionally see that the episode total cost $G_{c}(\tau_{k})$ decreases as $k$ increases, ideally below some user-given threshold, say $\delta$. Moreover, when the user is more tolerant of the cost (i.e., with a larger $\delta$), the agent should obtain higher rewards (i.e., larger $G(\tau_{k})$). We formalize these desiderata as the following constrained problem:
$\textstyle\max_{\theta}\mathbb{E}_{\pi_{\theta}}[\sum_{k=1}^{K}G(\tau_{k})]\mbox{\quad s.t.\quad}\forall k,\mathbb{E}_{\pi_{\theta}}\quantity[G_{c}(\tau_{k})]\leq\delta.$ (4)
Here, $K$ is the number of episodes that are allowed in the test time. $\tau_{k}$ is the $k$-th epsiode sampled from the policy $\pi_{\theta}$. The policy $\pi_{\theta}$ can depends on the context (i.e., $\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k})$) but the neural network parameters $\theta$ is fixed for all the $K$ episodes. We note that (4) is defined on a single test MDP that the algorithm has no access to during pretraining. This test MDP is usually sufficiently different from the MDPs that the algorithm has access to during pretraining. As a result, the agent cannot solve this test MDP via memorizing some good policies in the pretraining MDPs. Instead, it has to perform reinforcement learning over the $K$ episodes, although the parameters $\theta$ are fixed. Notably, although we formulate our problem in MDPs, the tasks we work on are only partially observable. This is a widely used practice in the RL community to promote the clarity of presentation (Mnih et al., 2015). We now explore different pretraining methods for solving the above problem.
Safe Supervised Pretraining.
We first establish safe supervised pretraining as a baseline approach. Similar to naive supervised pretraining, we collect a dataset $\quantity{\Xi_{i}}$ by executing various safe RL algorithms (e.g., Tessler et al. (2018); Ray et al. (2019)) in various safety-sensitive tasks (e.g., Ji et al. (2023); Gu et al. (2025)). Now each trajectory $\Xi_{i}$ consists of multiple episodes and each episode contains both the reward and the cost (cf. (3)). To further guide the action generation, we additionally condition the policy on both the return-to-go (RTG) $G_{t}(\tau)$ and the cost-to-go (CTG) $G_{c,t}(\tau)$. RTG is the total future rewards in the current episode, i.e., $G_{t}(\tau)\doteq\sum_{i=t+1}^{T}R_{t}$, and CTG is the total future cost in the current episode, i.e., $G_{c,t}(\tau)\doteq\sum_{i=t+1}^{T}C_{t}$. The RTG and CTG are inspired by Liu et al. (2023) for constrained decision Transformers and are designed to bias the agent towards achieving the specified return and cost. The main difference from Liu et al. (2023) is that in addition to RTG and CTG, we use entire episodes in input, while Liu et al. (2023) only includes states and actions but do not include per-step reward and per-step cost. As a result, Liu et al. (2023) still misses the opportunity to learn the algorithm demonstrated in the dataset, and they demonstrated only limited generalization capability. Specifically, given a trajectory $\Xi=(\tau_{1},\dots,\tau_{K})$ in the dataset, the safe supervised pretraining loss at a state $S_{t}^{k}$ is
$-\log\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})).$ (5)
We recall that the context $H_{t}^{k}$ is defined as $(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},C_{1}^{k},\dots,S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k},C_{t}^{k})$. In the pretraining, we do have access to RTG and CTG since the episode is already complete. But in the test time, we do not have access to them since they are calculated based on future rewards and costs, which are not available yet. Instead, we replace them with a spectrum of target RTG and CTG values. This gives the user the opportunity to control the trade-off between rewards and costs in the test time without fine-tuning the parameters. For example, in the test time, if the ratio RTG / CTG is small, it means the user is more risk-tolerant. If the ratio RTG / CTG is large (in the extreme case, CTG = 0), it means the user is more safety-focused.
Although the proposed safe supervised pretraining approach is to our knowledge novel, it is indeed a straightforward extension of existing unconstrained superivsed pretraining approaches like Laskin et al. (2023). We regard this as a baseline approach for sanity check. Any meaningful safe ICRL method should at least outperform this baseline.
Safe Reinforcement Pretraining.
Our safe supervised learning follows the offline pretraining paradigm. It is well-known that the performance of offline trained policies is usually worse than the online trained ones when online training is plausible (Vinyals et al., 2019; Mathieu et al., 2021). Motivated by this, we now study safe reinforcement pretraining. We note that even when we consider online pretraining, we still do not have access to the test MDP where (4) is defined. By online pretraining, we mean that we can interact with the pretraining MDPs directly. By contrast, offline supervised pretraining has access to only datasets collected from pretraining MDPs.
The underlying hypothesis of reinforcement pretraining is that by minimizing the loss (2) in a wide range of pretraining MDPs, the sequence model should be able to discover its own RL algorithm and implement it in the forward pass (see Lu et al. (2023); Grigsby et al. (2024a; b) for empirical evidence and Wang et al. (2025b) for theoretical evidence). Following this hypothesis, we now design an algorithm to solve (4) on a single pretraining MDP. To solve (4), primal-dual approaches in safe-RL (Achiam et al., 2017; Tessler et al., 2018; Liang et al., 2018; Ray et al., 2019) convert it to the dual problem below [^1] We omit $\theta$ and use $\max_{\pi}f(\pi)$ instead of $\max_{\theta}f(\pi_{\theta})$ for simplicity.:
$\textstyle\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})=\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}\quantity[{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\sum_{k=1}^{K}\lambda_{k}({\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta)]$ (6)
where ${\bm{\lambda}}\succeq 0$ is a componentwise inequality for the Lagrangian multipliers. Notably, the function class of the policy $\pi$ in $\max_{\pi}$ is all the policies in the form of (5) but without RTG, i.e., the policy can depend on the context and the CTG.
However, we observe that while the CMDP formulation requires a separate constraint and Lagrangian multiplier for each evaluation episode $\tau_{k}$, this is not desirable in our problem, as it would need to fix the number of evaluation episodes at the beginning of pretraining. Using fewer episodes would result in some multipliers being updated fewer times than others, which makes the optimization unsable and using more episodes is not possible.
Furthermore, there should be symmetry among all the constraints since they are all coming from a single cost function but simply correspond to different episodes. These encourage the use of a single multiplier for all episodes. However, when using a single multiplier, we would need a new way to not over-penalize the policy in the episodes that are already satisfying the constraint. To this end, We propose a modified Lagrangian function and an iterative optimization scheme and show it is more stable empirically while satisfying the desirable optimization properties. Let $g_{k}(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta$. We consider the following surrogate objective for the policy
$\textstyle L_{\Sigma}(\pi,\lambda)={\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\lambda\sum_{k=1}^{K}[g_{k}(\pi)]_{+}$
where $x]_{+}\doteq\max\quantity{x,0}$ and perform the following iterative updates:
$\textstyle\pi_{t+1}\in\operatorname*{argmax}_{\pi}L_{\Sigma}(\pi,\lambda_{t})\,\quad\lambda_{t+1}=[\lambda_{t}+\eta\,\max_{k}g_{k}(\pi_{t+1})]_{+}\quad\eta>0.$ (7)
Assumption 1.
The expected return $\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]$} and expected cost $\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]$} are bounded and continuous in $\pi$. The constrained problem (4) admits an optimal feasible policy $\pi^{\star}$.
Assumption 2.
[Paternain (2018)] Problem (4) satisfies conditions ensuring zero duality gap with (6) and there exists optimal Lagrange multipliers ${\bm{\lambda}}^{\star}\succeq 0$ such that
$\textstyle\underset{\text{s.t.}\ \forall k,{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}\leq\delta}{\max_{\pi}}{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}\ \;=\;\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})$
We now show that the set of the fixed points of (7) and the set of the optimizers of (4) are equal.
Theorem 1.
[Proof in Appendix B.1] We say a pair $(\bar{\pi},\bar{\lambda})$ is a fixed point of (7) if $\bar{\pi}\in\arg\max_{\pi}L_{\Sigma}(\pi,\bar{\lambda})$, and for all sufficiently small $\eta>0$, $\bar{\lambda}=[\bar{\lambda}+\eta\,\max_{k}g_{k}(\bar{\pi})]_{+}$. Let Assumptions 1 - 2 hold. Then every fixed point $(\bar{\pi},\bar{\lambda})$ of (7) is feasible and primal-optimal for (4). Conversely, every primal-optimal policy $\pi^{*}$ admits a corresponding fixed point $(\pi^{*},\bar{\lambda})$ for any multiplier $\bar{\lambda}\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$, where $\bm{\lambda}$$^{\star}$ is an optimal dual solution.
To implement the policy update (7), we consider an actor-critic approach following Grigsby et al. (2024a). The pseudocode is provided in Algorithm 1. Since our update (7) resembles an exact penalty convex optimization technique when $\lambda\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$ (Nocedal & Wright, 2006), we call the resulting actor-critic algorithm Exact Penalty Policy Optimization (EPPO).
4 Experiments
We now empirically investigate the proposed safe ICRL algorithms. Specifically, we aim to answer the following two questions.
  • (i)
    Do the proposed safe ICRL algorithms demonstrate generalization to out-of-distribution (OOD) tasks in complex environments with safety constraints?
  • (ii)
    How effectively do the proposed safe in-context RL algorithms achieve flexible reward-cost tradeoffs under varying cost tolerance?
We evaluate our approach on two constrained environments: (1) SafeDarkRoom and (2) SafeDarkMujoco: Point and Car.
These are modified from DarkRoom (Laskin et al., 2023) and SafetyGym benchmark (Ji et al., 2023) respectively. SafeDarkRoom is a grid-world setting where the agent can only perceive its own position and lacks visibility of the goal or obstacle locations. Rewards are sparse, with a positive reward obtained upon reaching the single goal in the map. Costs arise from multiple obstacles scattered across the environment. For instance, 25 obstacles in a $\times9$ grid map, which incur a cost of 1 each time the agent steps on one of the obstacles. To succeed, the agent must learn obstacle positions based on encountered cost signals, introducing additional complexity compared to goal-only environments due to the presence of multiple hazards. SafeDarkMujoco environment operates in continuous space with MuJoCo simulation (Todorov et al., 2012) but with a setup analogous to SafeDarkRoom. The robot senses its internal physical states, such as velocity, acceleration, position, and rotation angle, while all lidar inputs are turned off, preventing detection of obstacles and the goal. Consequently, the agent must develop exploratory behaviors to identify obstacles and goals, akin to SafeDarkRoom design, while navigating to the goal and minimizing collisions by learning from the previous collisions in the context.
Measuring OOD generalization.
Goal discovery-oriented environments such as DarkRoom are widely used in previous ICRL works (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025). In those works, the goals are randomly spawned over the map. Each goal corresponds to a new task. By ensuring the goals used in pretraining do not overlap with the goals used in testing, they ensure the test task is unseen during pretraining and can thus evaluate the generalization capability of their ICRL agents. However, such generalization can be achieved by interpolation (Kirk et al., 2021). For example, if the goals are spawned only on the black squares of a chessboard during pretraining, then an unseen goal location on one of the white squares can be viewed as an interpolation of the goals in the pretraining tasks. The agent thus may navigate to this unseen test goal by interpolating policies learned from pretraining. In other words, although those evaluation tasks are unseen during pretraining, it is not clear whether those tasks can fully demonstrate the challenges of OOD generalization.
To measure the OOD generalization of our proposed safe ICRL methods, we employ a challenging distance-based obstacle and goal spawning strategy: center-oriented for pretraining and edge-oriented for evaluation. The agent spawns at the map center, with obstacles and the goal distributed proportionally closer to this center during pretraining. During evaluation, obstacles and the goal follow an edge-oriented distribution. This approach applies similarly to the SafeDarkRoom and SafeNavigation environments. We argue that this setup is more challenging than those in prior studies (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025), as the agent must learn extrapolation rather than interpolation. More precisely, during pretraining, the grid position $(i,j)$ has an obstacle with probability $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$, where $c$ is the map center, $d(\cdot,\cdot)$ denotes Euclidean distance, and $\alpha>0$ to promote central density.
To generate an evaluation task, the grid position $(i,j)$ has an obstacle with probability $\mathbb{P}_{\text{test}}((i,j))\propto e^{\alpha d((i,j),c)}$ with $\alpha>0$ to favor edge density. The goals are generated similarly. This distributional shift is demonstrably OOD, both visually (Appendix C) and mathematically.
Proposition 1 (Proof in Appendix B.2).
The total variation distance satisfies $\lim_{\alpha\rightarrow\infty}\delta(\mathbb{P}_{\text{train}},\mathbb{P}_{\text{test}})=\lim_{\alpha\rightarrow\infty}\frac{1}{2}\sum_{O}|\mathbb{P}_{\text{train }}(O)-\mathbb{P}_{\text{test }}(O)|=1$ (maximum separation). Similarly, the KL divergence holds $\lim_{\alpha\rightarrow\infty}D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)=\infty$.
Question (i).
We now demonstrate emergent safe learning behaviors in OOD test tasks, thus giving an affirmative answer to Question (i). We use supervised pretraining and reinforcement pretraining in Section 3 on pretraining tasks with center-oriented obstacles and goals. Particularly, for supervised pretraining, the source algorithm to generate the dataset is PPO-Lagragian (Ray et al., 2019). goals. For reinforcement pretraining, we update the model for 30,000 steps on SafeDarkRoom and 10,000 steps on SafeDarkMujoco using Algorithm 1.
After pretraining, we evaluate the learned agents in evaluation tasks with edge-oriented obstacles and goals. The agent is able to interact with the evaluation task for multiple episodes, but cannot update its network parameters. During evaluation, for supervised pretraining, we set RTG to 1.0 and CTG to 0.0, targeting successful performance while minimizing cost violations. Reinforcement pretraining is solely conditioned on CTG. To evaluate robustness to CTG values during evaluation, we set CTG to a value uniformly sampled from the interval $[1,15]$ for SafeDarkRoom and $[10,50]$ for SafeDarkMujoco. Precisely, we sample $100$ distinct CTG values and report the average performance over these $100$ distinct CTG values. We recall that although the safe supervised pretraining method we propose is novel, we mostly regard it as a straightforward baseline. As can be seen in Figure 1, safe supervised pretraining succeeds in SafeDarkRoom as the episode return grows and the episode cost decreases. This aligns with the previously observed success of algorithm distillation in DarkRoom (Laskin et al., 2023; Zisman et al., 2023). But this baseline approach fails in SafeDarkMujoco. We conjecture that this is because of insufficiency in the dataset, as supported by ablation studies (Figure 6). By contrast, our safe reinforcement pretraining approach consistently succeeds in all three tested domains. It is also worth highlighting that in SafeDarkRoom, the safe reinforcement pretraining enables much faster adaptation to the new evaluation task than the safe supervised pretraining approach.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
Figure 1: Evaluation performance of superivsed and reinforcement pretraining. The curves are averaged over 1616 distinct OOD evaluation tasks, with each task featuring edge-oriented goal and obstacles. The shaded regions indicate standard errors. The $x$-axis is the episode index $k$ during the evaluation task. The $y$-axis is the corresponding episode return $G(\tau_{k})$ and the epsiode cost $G_{c}(\tau_{k})$ respectively. The straightforward safe supervised pretraining baseline succeeds only in SafeDarkRoom, while our novel safe reinforcement pretraining method succeeds in all three domains.
Question (ii).
We now demonstrate that we can adjust the behavior of the pretrained policy in test time by simply changing the RTG and/or CTG without making any parameter updates. Intuitively, with a higher CTG (i.e., a higher cost tolerance), the agent should afford bolder exploration and thus obtain higher rewards. Conversely, a lower CTG should promote conservative policies that prioritize safety, potentially at the expense of suboptimal rewards.
For the supervised pretrained model, we initialize RTG at 0.5 and CTG at 0.0, then adjust RTG = $\max\{0.5,\frac{\text{CTG}}{10}\}$ as CTG increases from 5 to 15 for SafeDarkRoom and 0 to 10 for SafeDarkMujoco. This setup enables gradual increases in both RTG and CTG to demonstrate balanced reward-cost trade-offs, consistently applied to SafeDarkRoom and SafeDarkMujoco. However, the supervised pretrained model fails to learn trade-off behaviors, exhibiting random patterns with varying RTG and CTG (Figure 2.(b)). We hypothesize that this stems from insufficient CTG diversity during training, leading to dataset inadequacy.
In our reinforcement pretrained model, we focus on controlling the CTG exclusively. This design simplifies the regulation of RTG based on the CTG. Simultaneously controlling both RTG and CTG is challenging, as determining the feasibility of RTG/CTG pairs is complex (Liu et al., 2023). Instead, the reinforcement pretrained model is designed to control CTG and pursue the maximal possible return for a given CTG, eliminating the need to justify the feasibility of specific solutions. Figure 2.(a) illustrates this relationship, showing that as cost limits increase, the return also increases.
Supervised pretraining depends on a long replay buffer to learn from logged learning histories. Due to context length constraints, we need to subsample trajectories from these learning histories for supervised pretraining. These histories are sampled from distinct learning phases: early (low return, high cost), middle (average return, average cost), and expert (high return, low cost). To make sequence modeling feasible, trajectories need to be labeled with RTG, which provides information about the expected phase (Dai et al., 2024). However, this approach requires significantly more data, as noted by Liu et al. (2023). In contrast, reinforcement pretraining leverages online interaction histories, eliminating the need to distinguish between different data phases.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
Figure 2: Evaluation performance of supervised and reinforcement pretraining with varying CTG. The curves are displayed across a range of cost limits. For each cost limit, the result is averaged over 1616 distinct OOD evaluation tasks, with each task incorporating edge-oriented goals and obstacles. The shaded regions indicate standard errors. The $x$-axis is the CTG. The $y$-axis is the total episode return (i.e., $\sum_{k=1}^{K}G(\tau_{k})$) and the maximum episode cost (i.e., $\max_{k\in K}G_{c}(\tau_{k})$) with the corresponding CTG. The policy from safe reinforcement pretraining succeeds in converting a higher cost limit (i.e., a higher CTG) to a higher return while the policy from safe supervised pretraining fails to do so.
Ablations.
We further conduct ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, alongside specific factors such as the dataset size for supervised pretraining. Our findings show distinct sensitivities to these factors. Reinforcement pretraining is largely unaffected by model size, whereas model size significantly influences supervised pretraining performance. For context length, reinforcement pretraining performs better with longer sequences, while supervised pretraining performs worse. Conversely, supervised pretraining excels with shorter context lengths, while reinforcement pretraining performs poorly. These results suggest that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, handles long-term credit assignment more effectively. We confirm the well-established finding that supervised pretraining is highly sensitive to dataset size. When the dataset is small, the model fails to learn, exhibiting random behavior on out-of-distribution data. In Figure 5.(c), we compare EPPO to a naive primal–dual solver for (6) with per-episode multipliers $\lambda$. The overall structure of this naive algorithm resembles RCPO (Tessler et al., 2018). EPPO trains more stably with minimal tuning and solves SafeDarkRoom and SafeDarkMujoco without changes beyond time-limit related settings (Table 1). The detailed evaluation curves of the ablation study are provided in Figures 56 in Appendix E. We evaluate a variant of safe supervised pretraining incorporating algorithm distillation with noise in Appendix D. The results indicate that learning histories generated by noisy actions do not effectively represent true learning histories in complex environments with safety constraints.
5 Related Works
ICRL.
Learning to improve on a new MDP by interacting with it and without parameter updates is first studied in meta RL (Duan et al., 2016; Wang et al., 2016). See Beck et al. (2025) for a detailed survey of meta RL. In general, the demonstrated zero parameter update generalization is weak in meta RL literature. Most of them rely on task identification, i.e., identifying pretraining tasks that are similar to the evaluation task and acting as if the evaluation task were the pretraining tasks. Laskin et al. (2023) coined the word ICRL and demonstrated generalization to evaluation tasks that are far away from the pretraining tasks, which spurs increasing attention in ICRL. Variations of algorithm distillation have been proposed for efficient in-context reinforcement learning, including the Decision Pretrained Transformer (Lee et al., 2023) and Algorithm Distillation with Noise (Zisman et al., 2023). Both approaches aim to efficiently train transformer models for in-context learning using optimal policies. See Moeini et al. (2025) for a detailed survey of ICRL. The main novelty of this work is to first study safe ICRL within the constrained MDP framework.
Safe RL.
Safe RL methods commonly adopt CMDP formulations to enforce compliance with safety constraints during exploration and policy optimization (Garcıa & Fernández, 2015; Gu et al., 2022; Wachi et al., 2024). Constrained Policy Optimization (CPO) serves as a foundational algorithm in this area, effectively balancing performance rewards with safety requirements (Achiam et al., 2017; Wachi & Sui, 2020). Subsequent extensions, such as Constrained Update Projection and Constrained Reward Policy Optimization, have built upon CPO to enhance theoretical guarantees and practical applicability in safe RL scenarios (Xu et al., 2021; Yang et al., 2022). Shielding with function encoders and conformal prediction has been proposed to handle unseen OOD environments (Kwon et al., 2025). However, this work emphasizes runtime adaptation rather than learning algorithms through contextual information. Reward-Constrained Policy Optimization (Tessler et al., 2018) is the closest work to EPPO, but it enforces safety only in expectation rather than at the episode level. From the offline RL perspective, the Constrained Decision Transformer (Liu et al., 2023) leverages the transformer architecture for safe RL. However, existing approaches enforce safety constraints only when the evaluation task closely matches the pretraining tasks. In contrast, our safe ICRL framework verifies safety constraints even on evaluation tasks that differ substantially from the pretraining distribution.
Safe Meta RL.
Safe meta RL has emerged to enable fast adaptation to new tasks while enforcing safety constraints. Earlier work (Luo et al., 2021) uses a three-phase approach that meta-learns a safety critic from offline data across multiple environments, adapts it to a new environment with a smaller offline dataset, and employs the adapted safety critic with a recovery policy to ensure safe learning while minimizing constraint violations. Khattar et al. (2023) provides task-averaged regret bounds for rewards and constraints via gradient-based meta-learning. More recently, Guan et al. (2024) proposes a cost-aware context encoder that uses supervised cost relabeling and contrastive learning to infer tasks based on safety constraints. To our knowledge, all previous safe meta RL works require parameter updates to achieve safe adaptation in evaluation tasks. By contrast, our work on safe ICRL enables safe adaptation to evaluation tasks without parameter updates.
6 Conclusion
This work pioneers the study of safe ICRL, examining both reinforcement pretraining and supervised pretraining. We develope a safe supervised pretraining appraoch based on the constrained decision transformer, showcasing its capability for OOD generalization while revealing its dependence on dataset size. We introduce EPPO, a safe reinforcement pretraining approach via exact penalty policy optimization to enforce strict cost constraints within each episode. Theoretically, we prove that EPPO’s fixed point satisfies strict safety constraints while maximizing rewards. We conduct a thorough empirical study of the proposed approaches in challenging benchmarks that provably demand OOD generalization (Proposition 1). Such demand for OOD generalization is missing in many benchmarks used by previous works (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025) Through those challenging benchmarks, we confirm that our pretrained safe ICRL agents adapt to not only unseen but also OOD evaluation tasks while respecting the safety constraints. We envision this work guiding the development of robust, generalizable, and safe RL algorithms for real-world applications.
Acknowledgments
This work is supported in part by the US National Science Foundation under the awards III-2128019, SLES-2331904, CAREER-2442098, and CCF-1942836, the Commonwealth Cyber Initiative’s Central Virginia Node under the award VV-1Q26-001, and an Nvidia academic grant program award.
Appendix A Algorithm
1:Input: discount factor $\gamma$, cost threshold $\delta$, number of training steps $T_{\max}$, batch size NN, env time limit $t_{\max}$, episodes-per-history range $K_{\min},K_{\max}]$, CTG range $\mathrm{CTG}_{\min},\mathrm{CTG}_{\max}]$, environment distribution $\mathcal{E}$
2:Initialize: actor $\pi(\cdot\mid s,H,\mathrm{CTG};\theta_{p})$, reward critic $Q(s,a;\theta_{v})$, cost critic $Q^{c}(s,a;\theta_{c})$, and target nets $\{\pi^{\prime},Q^{\prime},Q^{c\prime}\}$. $\lambda\leftarrow 0$,  replay buffer $R\leftarrow\varnothing$
3:for $T\leftarrow 1$ to $T_{\max}$ do
4:  $K\leftarrow\texttt{rnd}(K_{\min},K_{\max})$,  $\mathrm{CTG}\leftarrow\texttt{rnd}(\mathrm{CTG}_{\min},\mathrm{CTG}_{\max})$,  $t\leftarrow 0$
$H\leftarrow[\;]$ // list of episodes
[…]
6:  for $k\leftarrow 1$ to $K$ do
7:   $t_{\text{start}}\leftarrow t$,  $\mathrm{CTG}_{k}\leftarrow\mathrm{CTG}$,  $e_{k}\leftarrow[\;]$ // list of transitions
8:   while $t-t_{\text{start}}<t_{\max}$ and $s_{t}$ not terminal do
9:    sample $a_{t}\sim\pi_{\theta_{p}}(\,\cdot\mid s_{t},H,\mathrm{CTG}_{k})$
10:    step $a_{t}$ in env $\to(s_{t+1},r_{t+1},c_{t+1})$
11:    append $(s_{t},a_{t},r_{t+1},c_{t+1},s_{t+1},\mathrm{CTG}_{k})$ to $e_{k}$
12:    $\mathrm{CTG}_{k}\leftarrow\mathrm{CTG}_{k}-c_{t+1}$,  $t\leftarrow t+1$
13:   end while
14:   append $e_{k}$ to $H$
15:   $s_{t}\leftarrow\texttt{env.reset()}$
16:  end for
17:  append $H$ to $R$
18:  $\mathcal{D}\leftarrow\texttt{Sample}(R,N)$ // NN trajectories
19:  reset accumulators: $d\theta_{v}\leftarrow 0$, $d\theta_{c}\leftarrow 0$, $d\theta_{p}\leftarrow 0,d\lambda\leftarrow 0$
20:  for each trajectory $H\in\mathcal{D}$ do
21:   for each episode $e\in H$ do
22:    $C_{e}\leftarrow\sum_{c\in e}c$,  $C_{e}^{\max}\leftarrow$-$\infty$,  $v_{e}\leftarrow\mathbf{1}\{C_{e}>\delta\}$
23:    for each $(s_{t},a_{t},r_{t+1},c_{t+1},s_{t+1},\mathrm{CTG}_{t})\in e$ do
24:     sample $a^{\prime}_{t+1}\sim\pi_{\theta_{p}^{\prime}}(\,\cdot\mid s_{t+1},H_{\leq t+1},\mathrm{CTG}_{t}-c_{t+1})$
25:     $L_{v}\leftarrow\big(Q_{\theta_{v}}(s_{t},a_{t})-[\,r_{t+1}+\gamma Q_{\theta_{v}^{\prime}}(s_{t+1},a^{\prime}_{t+1})\,]\big)^{2}$
26:     $d\theta_{v}\leftarrow d\theta_{v}+\nabla_{\theta_{v}}L_{v}$
27:     $L_{c}\leftarrow\big(Q^{c}_{\theta_{c}}(s_{t},a_{t})-[\,c_{t+1}+\gamma Q^{c}_{\theta_{c}^{\prime}}(s_{t+1},a^{\prime}_{t+1})\,]\big)^{2}$
28:     $d\theta_{c}\leftarrow d\theta_{c}+\nabla_{\theta_{c}}L_{c}$
29:     sample $\hat{a}_{t}\sim\pi_{\theta_{p}}(\,\cdot\mid s_{t},H_{\leq t},\mathrm{CTG}_{t})$
30:     $L_{p}\leftarrow-\,Q_{\theta_{v}}(s_{t},\hat{a}_{t})+\lambda\,v_{e}\,Q^{c}_{\theta_{c}}(s_{t},\hat{a}_{t})$
31:     $d\theta_{p}\leftarrow d\theta_{p}+\nabla_{\theta_{p}}L_{p}$
32:    end for
33:    $C_{e}^{\max}\leftarrow\max\quantity{C_{e}^{\max},C_{e}}$
34:   end for
35:   $d\lambda\leftarrow d\lambda+C_{e}^{\max}$
36:  end for
37:  apply parameter updates to $\theta_{v},\theta_{c},\theta_{p},\lambda$ using $d\theta_{v},d\theta_{c},d\theta_{p},d\lambda$
38:  update targets $\pi^{\prime},Q^{\prime},Q^{c\prime}$
39:end for
Algorithm 1 EPPO Implemented with DDPG
Note that our proposed optimization (7) can be used with any policy optimizer. In our implementation, we chose DDPG following Grigsby et al. (2024a) to support continuous and discrete action spaces and to enable highly parallel data-collection while doing off-policy training.
Appendix B Proofs
We organize the proofs into two subsections: one for Theorem 1 and one for Proposition 1. For Theorem 1, we first prove the supporting lemma and then Theorem 1 itself. For Proposition 1, we provide proofs for both Total Variation Distance and KL Divergence.
B.1 Proof of Theorem 1
Lemma 1.
[…]
Proof.
Consider two cases.
Case $\bar{\lambda}>0$. For all small $\eta>0$, $\bar{\lambda}+\eta\bar{s}>0$, hence
$\bar{\lambda}+\eta\bar{s}]_{+}=\bar{\lambda}+\eta\bar{s}.$
The equality $\bar{\lambda}=[\bar{\lambda}+\eta\bar{s}]_{+}$ then forces $\bar{\lambda}=\bar{\lambda}+\eta\bar{s}$, so $\bar{s}=0$. Thus $\bar{s}\leq 0$ and $\bar{\lambda}\bar{s}=0$ hold.
Case $\bar{\lambda}=0$. Then we have $0=[\eta\bar{s}]_{+}=\max\{0,\eta\bar{s}\}$ for all small $\eta>0$, which implies $\eta\bar{s}\leq 0$, hence $\bar{s}\leq 0$. Trivially $\bar{\lambda}\bar{s}=0$ and $\bar{\lambda}\geq 0$. ∎
We now proceed to the proof of Theorem 1.
Proof.
Let $J(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}$ in this proof. We prove the theorem in both directions. Also let $g_{k}^{+}(\pi)\coloneqq\max\{0,g_{k}(\pi)\}$, and $s(\pi)\coloneqq\max_{k}g_{k}(\pi)$.
1. Fixed point $\Rightarrow$ Primal optimal.
Feasibility. By the lemma 1, $s(\bar{\pi})\leq 0$. Since $s(\bar{\pi})=\max_{i}g_{i}(\bar{\pi})$, each $g_{i}(\bar{\pi})\leq 0$. Thus $\bar{\pi}$ is feasible.
Optimality. Let $\pi^{\star}$ be any optimal feasible policy (exists by assumption 1). On the feasible set, $g_{i}^{+}(\pi^{*})=0$, so $L_{\Sigma}(\pi^{*},\bar{\lambda})=J(\pi^{*})$. Because $\bar{\pi}$ maximizes $L_{\Sigma}(\cdot,\bar{\lambda})$,
$J(\bar{\pi})=L_{\Sigma}(\bar{\pi},\bar{\lambda})\ \geq\ L_{\Sigma}(\pi^{\star},\bar{\lambda})=J(\pi^{\star}).$
Conversely, by optimality of $\pi^{\star}$ among feasible policies and feasibility of $\bar{\pi}$, $J(\bar{\pi})\leq J(\pi^{\star})$. Hence $J(\bar{\pi})=J(\pi^{\star})$.
2. Primal optimal $\Rightarrow$ Fixed point.
Let $\lambda^{\star}$ be an optimal dual solution (assumption 2). Then by strong duality,
$p^{\star}\;=\;\max_{\pi}\Big(J(\pi)-\textstyle\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\Big).$
Therefore, for every $\pi$,
$J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\ \leq\ p^{\star}.$ (8)
Using $\lambda_{i}^{\star}\leq\|\lambda^{\star}\|_{\infty}$ and $g_{i}^{+}(\pi)\geq g_{i}(\pi)$,
$\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\ \leq\ \|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi).$
Subtract this from $J(\pi)$ and combine with (8):
$L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big)\;=\;J(\pi)-\|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi)\;\leq\;J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\;\leq\;p^{\star}.$ (9)
For a primal-optimal $\pi^{\star}$, feasibility gives $\sum_{i}g_{i}^{+}(\pi^{\star})=0$, hence
$L_{\Sigma}\big(\pi^{\star},\ \|\lambda^{\star}\|_{\infty}\big)=J(\pi^{\star})=p^{\star}.$
Together with (9), this shows
$\pi^{\star}\in\arg\max_{\pi}L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big).$
It remains to check multiplier stationarity at $\bar{\lambda}=\|\lambda^{\star}\|_{\infty}$. By complementary slackness, $\sum_{i}\lambda_{i}^{\star}g_{i}(\pi^{\star})=0$. If $\|\lambda^{\star}\|_{\infty}>0$, at least one constraint is active at $\pi^{\star}$, so $\max_{i}g_{i}(\pi^{\star})=0$. Hence $\bar{\lambda}+\eta\,s(\pi^{\star})]_{+}=[\bar{\lambda}+\eta\cdot 0]_{+}=\bar{\lambda}$ for all small $\eta>0$. If $\|\lambda^{\star}\|_{\infty}=0$, $\pi^{\star}$ is unconstrained-optimal with $s(\pi^{\star})\leq 0$, and $0+\eta\,s(\pi^{\star})]_{+}=0$. In both cases the projection is stationary. Thus $(\pi^{\star},\bar{\lambda})$ with $\bar{\lambda}=\|\lambda^{\star}\|_{\infty}$ is a fixed point.
On the multiplier update.
If $s(\pi_{t+1})>0$, then $\lambda_{t+1}=[\lambda_{t}+\eta s(\pi_{t+1})]_{+}\geq\lambda_{t}$. If constraints are strict ($s(\pi_{t+1})<0$), then $\lambda_{t+1}\leq\lambda_{t}$. Under persistent violation, $\lambda_{t}$ eventually exceeds $\|\lambda^{*}\|_{\infty}$, by the exact-penalty bound (9), any maximizer of $L_{\Sigma}(\cdot,\lambda_{t})$ is then feasible/optimal and $s(\pi_{t+1})\leq 0$, so $\lambda$ stabilizes. ∎
B.2 Proof of Proposition 1
Total Variation Distance
First, we present the exact form of probabillites for $\mathbb{P}_{\text{train}}(O)$ and $\mathbb{P}_{\text{test}}(O)$ where $O\in\mathcal{O}=\{(i,j)\}_{1\leq i\leq n,1\leq j\leq m}$. Let the map of center $c=(i_{c},j_{c})$. Then, $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$ and $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$ implies:
$\mathbb{P}_{\text{train}}((i,j))=\frac{e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}}\quad\text{and}\quad\mathbb{P}_{\text{test}}((i,j))=\frac{e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{\alpha}},$
[…]
$\delta(\mathbb{P}_{\text{train}},\mathbb{P}_{\text{test}})$ $=\frac{1}{2}\sum_{O\in\mathcal{O}}|\mathbb{P}_{\text{train}}(O)-\mathbb{P}_{\text{test}}(O)|$ (10)
$=\frac{1}{2}\sum_{O\in\mathcal{O}}(|\mathbb{P}_{\text{train}}(O)|+|\mathbb{P}_{\text{test}}(O)|)=1.$ (11)
Now, we claim that $supp(\mathbb{P}_{\text{train}})\cap supp(\mathbb{P}_{\text{test}})=\emptyset$ as $\alpha\rightarrow\infty$. For the training distribution $\mathbb{P}_{\text{train}}$, the term $e^{-\alpha\left(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}\right)}\rightarrow 0$ if $\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}>0$. Hence,
$Z_{-\alpha}=e^{-\alpha\cdot 0}+\sum_{(i,j)\neq\left(i_{c},j_{c}\right)}e^{-\alpha\left(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}\right)}\rightarrow 1$
Hence, only when $i=i_{c}$ and $j=j_{c}$, we have positive probability $\mathbb{P}_{\text{train}}((i_{c},j_{c}))=\frac{e^{-\alpha\cdot 0}}{Z_{-\alpha}}=\frac{1}{Z_{-\alpha}}=1$. Thus $supp(\mathbb{P}_{\text{train}})\rightarrow\{(i_{c},j_{c})\}$
For the test distribution $\mathbb{P}_{\text{test}}$, the term $e^{\alpha\left(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}\right)}$ is maximized when $\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}$ is maximized. Let $d_{\max}=\max_{(i,j)\in\mathcal{O}}\left(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}\right)$, and $\mathcal{O}_{\text{edge }}=\left\{(i,j)\in\mathcal{O}\mid\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}=d_{\max}\right\}$. Then the partial function can be decomposed into two terms:
$Z_{\alpha}=\sum_{(i,j)\in\mathcal{O}_{\text{edge }}}e^{\alpha d_{\max}}+\sum_{(i,j)\notin\mathcal{O}_{\text{edge }}}e^{\alpha\left(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}\right)}.$
Hence, as $\alpha\rightarrow\infty$, $\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right)^{2}<d_{\text{max}}$ implies that $\frac{e^{\alpha(\left(i-i_{c}\right)^{2}+\left(j-j_{c}\right))}}{Z_{\alpha}}\rightarrow 0$. Thus, we obtain
$\mathbb{P}_{\text{test }}((i,j))\rightarrow\begin{cases}\frac{1}{\left|\mathcal{O}_{\text{edge }}\right|}&\text{ if }(i,j)\in\mathcal{O}_{\text{edge }},\\ 0&\text{ otherwise }\end{cases}\quad as\quad\alpha\rightarrow\infty.$
Hence, $supp(\mathbb{P}_{\text{test}})\rightarrow\mathcal{O}_{\text{edge}}.$ Finally, we conclude that $supp(\mathbb{P}_{\text{train}})\cap supp(\mathbb{P}_{\text{test}})=\{(i_{c},j_{c})\}\cap\mathcal{O}_{\text{edge}}=\emptyset$ as $\alpha\rightarrow\infty.$
KL Divergence
We keep the same notation to the proof so far. By the definition of KL divergence and explicit form of the probabilities, we have
$D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)$ $=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{\mathbb{P}_{\text{train }}((i,j))}{\mathbb{P}_{\text{test }}((i,j))}$ (12)
$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{Z_{\alpha}e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}$ (13)
$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\left(\log\frac{Z_{\alpha}}{Z_{-\alpha}}-2\alpha((i-i_{c})^{2}+(j-j_{c})^{2})\right).$ (14)
As we have shown in the previous proof for TV distance, $Z_{\alpha}\rightarrow|\mathcal{O}_{\text{edge}}|e^{\alpha d_{\text{max}}}$ and $Z_{-\alpha}\rightarrow 1$. Hence, for pairs $(i,j)\not=(i_{c},j_{c})$, the associated term $\mathbb{P}_{\text{train }}((i,j))\left(\log\frac{Z_{\alpha}}{Z_{-\alpha}}-2\alpha((i-i_{c})^{2}+(j-j_{c})^{2})\right)$ goes to $\mathbb{P}_{\text{train }}((i,j))\left(\log|\mathcal{O}_{\text{edge}}|+\alpha d_{\text{max}}-2\alpha d((i,j),c)\right)$. For $(i,j)\not=(i_{c},j_{c})$ and $d_{\text{min}}=\min_{(i,j)\neq\left(i_{c},j_{c}\right)}d((i,j),c)$, we have
$P_{\text{train }}((i,j))=\frac{e^{-\alpha d((i,j),c)}}{Z_{-\alpha}}\leq e^{-\alpha d_{\text{min}}}\rightarrow 0.$ (15)
Since the exponential rate converges to zero more rapidly than the linear rate with respect to $\alpha$, the term $\mathbb{P}_{\text{train }}((i,j))\left(\log|\mathcal{O}_{\text{edge}}|+\alpha d_{\text{max}}-2\alpha d((i,j),c)\right)$ goes to 0 as $\alpha\rightarrow\infty$. Therefore, we only consider the term when $(i,j)=(i_{c},j_{c})$. But this term comes down to $\mathbb{P}_{\text{train}}((i,j))\log\frac{Z_{\alpha}}{Z_{-\alpha}}$, which goes to $\infty$ as $\alpha\rightarrow\infty$. Thus $D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)\geq\infty.$
Appendix C Training Details
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Training Environments(b) Test EnvironmentsSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
Figure 3: During training, goals and obstacles are generated with a center-oriented approach, while during evaluation, they are edge-oriented. This applies consistently to both goals and obstacles. We set $\alpha=0.5$ for generating goals and obstacles. The red color denotes the robot, the green color represents the goal location, and obstacles are depicted in shades of blue.
Environments.
We introduce the details of the environments. The environments are visualized in Figure 3. Both environment uses $\alpha=0.5$ to generate goals and obstacles.
For SafeDarkRoom, We use a 9x9 Grid and give one reward upon reaching the goal and incur one cost when going on an obstacle cell. We terminate the episode whenever the agent reaches the goal, which results in a truly sparse reward, often referred to as DarkRoom Hard.
For SafeDarkMujoco, a continuous state and action space counterpart to SafeDarkRoom, the agent lacks lidar information and is blind to goal and obstacle locations. Instead, it perceives its own position and rotation matrix. A sparse reward is obtained when the agent reaches the goal, terminating the episode. To enhance runtime efficiency, we employ macro actions, compressing $n$ simulation steps into a single step. For example, with $n=5$, a policy action is repeated over five internal simulation steps, reducing the default 250 simulation steps to $\frac{250}{n}$. In our experiments, we set $n=5$.
Reinforcement Pretraining.
While running Algorithm 1 for reinforcement pretraining, we resample a new environment from the training distribution described in Section 4 every $K$ episodes. For each environment, a CTG is also sampled, ranging from $[1,15]$ for SafeDarkRoom and $[10,50]$ for SafeDarkMujoco. The remaining hyperparameters are provided in Table 1.
Our architecture follows Grigsby et al. (2024a). We employ an MLP time-step encoder that maps each tuple $(S_{t},A_{t},R_{t},C_{t})$ to an embedding, which is then fed into a transformer-based trajectory encoder. A prediction head outputs either the action distribution (for discrete actions) or the value (for continuous actions).
In practice, since the environments are generated randomly, it is possible that an environment doesn’t have a feasible solution, causing the agent to inevitably violate the cost threshold. This results in $\lambda$ growing too large in an attempt to mitigate these cases. To avoid this, we suggest capping $\lambda$ to a reasonably large value. Ignoring abnormally high-cost episodes is another solution worth exploring.
Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
$K_{\min}$, $K_{\max}$ 50, 50 20, 20 20, 20
Episode time limit $t_{\max}$ 30 75 75
Replay buffer capacity 100,000 100,000 100,000
Embedding Dim 64 64 64
Hidden Dim 64 64 64
Num Layers 4 4 4
Num Heads 8 8 8
Seq Len 1500 1500 1500
Attention Dropout 0 0 0
Residual Dropout 0 0 0
Embedding Dropout 5 5 5
Learning Rate 3e-4 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0 1.0
Batch Size 32 32 32
Num Updates 30k 10k 10k
Optimizer Adam Adam Adam
Table 1: Parameters for Safe Reinforcement Pretraining
Supervised Pretraining.
In supervised pretraining, we collect a dataset $\mathcal{D}=\quantity{\Xi_{i}}$ comprising multiple trajectories, where each trajectory $\Xi_{i}\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$ represents a sequence of episodes generated by running existing safe RL algorithms on various CMDPs. Each episode $\tau_{k}$ in a trajectory $\Xi_{i}$ comprises states, actions, rewards, and costs, with the episode return $G(\tau_{k})\doteq\sum_{t=1}^{T}R_{t}$ and cost-to-go $G_{c,t}(\tau_{k})\doteq\sum_{i=t+1}^{T}C_{i}$, expected to increase and decrease, respectively, with $k$ as the RL algorithm learns. The policy $\pi_{\theta}$ is trained autoregressively using an imitation learning loss to distill the behavior of RL algorithms demonstrated in the dataset, following the concept of algorithm distillation (Laskin et al., 2023). For discrete action spaces (e.g., SafeDarkRoom), we use a cross-entropy loss, defined as:
$\mathcal{L}_{\mathrm{CE}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[-\log\pi_{\theta}(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right],$ (16)
where $\pi_{\theta}$ is the categorical distribution over discrete actions. For continuous action spaces (e.g., SafeDarkMujoco), we use an L2 loss, where the transformer outputs an action mean $\mu_{t}=\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))$, minimizing:
$\mathcal{L}_{\mathrm{L2}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[\left\|A_{t}^{k}-\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right\|^{2}\right].$ (17)
These losses extend the imitation learning loss from Equation (5) by incorporating RTG and CTG conditioning, aligning with our goal of distilling RL algorithms into the policy’s forward pass while optimizing for safety constraints.
Dataset Collection. For supervised pretraining, we collect learning trajectories using reinforcement learning (RL) algorithms. As our base RL algorithm, we employ PPO-Lagrangian Schulman et al. (2017); Ray et al. (2019), which is designed to maximize rewards while enforcing safety constraints. These trajectories capture behaviors that learn to avoid obstacles.
To introduce variation in the learned behaviors, we vary the cost limits in PPO-Lag across multiple settings. For the SafeDarkRoom environment, we use three cost limits: 0, 2.5, and 5.0. For the SafeDarkMujoco environment, we similarly use three cost limits: 0, 5, and 10. For each cost limit, we collect 50,000 steps of learning history in both environments.
Hyper Parameters.
We report the hyperparameters used for reinforcement pretraining (Table 1) and supervised pretraining (Table 2). Reinforcement pretraining adopts the AMAGO framework (Grigsby et al., 2024a) integrated with our novel EPPO method. Supervised pretraining employs a constrained decision transformer as the backbone (Liu et al., 2023).
Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
Embedding Dim 64 64 64
Hidden Dim 512 256 256
Num Layers 8 8 8
Num Heads 8 8 8
Seq Len 100 300 200
Attention Dropout 0.5 0.5 0.5
Residual Dropout 0.1 0.1 0.1
Embedding Dropout 0.3 0.3 0.3
Learning Rate 3e-4 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0 1.0
Batch Size 512 128 128
Num Updates 300k 500k 500k
Optimizer Adam Adam Adam
Table 2: Parameters for Safe Supervised Pretraining
Appendix D Variation of Supervised Pretraining
Figure 4: Performance Comparison of Algorithm Distillation (AD) style supervised pretraining and Algorithm Distillation with Noise (AD-EPS) style supervised pretraining in SafeDarkRoom. AD-EPS fails to generalize in OOD environments.
In this section, we compare the performance of Algorithm Distillation (AD) (Laskin et al., 2023) and Algorithm Distillation with Noise (AD-EPS) (Zisman et al., 2023) in SafeDarkRoom. AD-EPS is designed to learn in-context RL algorithms using datasets generated from a perturbed optimal policy. This approach allows efficient generation of learning trajectories from a single optimal policy. However, our results in complex environments with cost signals reveal that AD-EPS relies on artificial trajectories. AD-EPS fails to learn effective in-context RL algorithms when relying solely on perturbed optimal policies, as the perturbations do not account for obstacle avoidance in SafeDarkRoom environment and instead introduce random action variations.
Appendix E Ablation Studies
In this section, we present our ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, as well as specific factors like dataset size for supervised pretraining. Our findings reveal distinct sensitivities to these factors. Reinforcement pretraining remains largely unaffected by model size, whereas supervised pretraining performance is significantly influenced by model size (See Figures 5.(b) and 6.(b)). Regarding context length, reinforcement pretraining benefits from longer sequences, while supervised pretraining exhibits degraded performance with longer contexts. Conversely, supervised pretraining performs better with shorter context lengths, where reinforcement pretraining struggles (See Figures 5.(a) and 6.(a)). These results indicate that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, manages long-term credit assignment more effectively. We also confirm the well-established finding that supervised pretraining is highly sensitive to dataset size, with models failing to learn and exhibiting random behavior on out-of-distribution data when the dataset is small (See Figure 6.(c)).
(a) Context Length
(b) Model Size
(c) Optimization
Figure 5: Ablations of Safe Reinforcement Pretraining on SafeDarkRoom. (a), (b) The evaluation is set up similarly to Question (i). For context length, we compare 150 and 3000 against the base value of 1500. For model size, we compare embedding dimensions 32 and 128 against the base of 64. (c) At each training step, the average total return of 50 episodes across 10 random test environments, and the average Q Value across 50 episodes of 250 random train enironments are plotted. EPPO is easier to tune and more stable to train than the naive primal-optimal method.
(a) Context Length
(b) Model Size
(c) Dataset Size
Figure 6: Ablations of Safe Supervised Pretraining on SafeDarkRoom. The evaluation is set up similarly to Question (i). For context length, we use a smaller base model and test three sequence lengths: 100, 500, and 1000. For dataset size, large refers to the full dataset, medium uses 50% of the dataset, and small uses 5% of the original dataset. Model size increases with the number of hidden layers: 2, 4, and 8, keeping other factors constant.

### v2

@@
 Abstract
-In-context reinforcement learning (ICRL) is an emerging RL paradigm where the agent, after some pretraining procedure, is able to adapt to out-of-distribution test tasks without any parameter updates.
-The agent achieves this by continually expanding the input (i.e., the context) to its policy neural networks.
-For example, the input could be all the history experience that the agent has access to until the current time step.
-The agent’s performance improves as the input grows, without any parameter updates.
-In this work, we propose the first method that promotes the safety of ICRL’s adaptation process in the framework of constrained Markov Decision Processes.
-In other words, during the parameter-update-free adaptation process, the agent not only maximizes the reward but also minimizes an additional cost function.
-We also demonstrate that our agent actively reacts to the threshold (i.e., budget) of the cost tolerance.
-With a higher cost budget, the agent behaves more aggressively, and with a lower cost budget, the agent behaves more conservatively.
+In-context reinforcement learning (ICRL) is an emerging RL paradigm where an agent, after pretraining, can adapt to out-of-distribution test tasks without any parameter updates, instead relying on an expanding context of interaction history.
+While ICRL has shown impressive generalization, safety during this adaptation process remains unexplored, limiting its applicability in real-world deployments where test-time behavior is expected to be safe.
+In this work, we propose SCARED: Safe Contextual Adaptive Reinforcement via Exact-penalty Dual, the first method that promotes safe adaptation of ICRL under the constrained Markov decision process framework.
+During the parameter-update-free adaptation process, our agent not only maximizes the reward but also keeps the accumulated cost within a user-specified safety budget.
+We also demonstrate that the agent actively reacts to the safety budget; with a higher safety budget, the agent behaves more aggressively, and with a lower safety budget the agent behaves more conservatively.
+Across challenging benchmarks, SCARED consistently enables safe and robust in-context adaptation, outperforming existing ICRL and safe meta-RL baselines.
+In-context Reinforcement Learning, Safe Reinforcement Learning
 1 Introduction
-Reinforcement learning (RL, Sutton & Barto (2018)) is a framework for solving sequential decision making problems via trial and error.
-Modern RL agents are usually parameterized by deep neural networks, and the trial-and-error process is typically achieved by updating the parameters of the neural networks (Mnih et al., 2015).
+Reinforcement learning (RL, Sutton & Barto (2018)) has achieved remarkable success in sequential decision-making, from game playing (Mnih et al., 2015) to robotic control (Levine et al., 2016).
+These achievements rely on deep neural networks whose parameters are updated through trial-and-error interaction with the environment.
+However, parameter updates at test time are not always feasible for adaptation: gradient computation may be unavailable, careful hyperparameter calibration can be required, and fine-tuning in meta-learning settings risks catastrophic forgetting of previously learned behaviors (Kirkpatrick et al., 2017).
 Recently, in-context reinforcement learning (ICRL, Moeini et al.
-(2025)) emerges as an RL paradigm that achieves this trial-and-error process without any neural network parameter updates (Duan et al., 2016; Mishra et al., 2018; Xu et al., 2022; Kirsch et al., 2023; Lin et al., 2023; Lu et al., 2023; Raparthy et al., 2023; Sinii et al., 2023; Grigsby et al., 2024a; b; Cook et al., 2024; Dai et al., 2024; Dong et al., 2024; Elawady et al., 2024; Huang et al., 2024; Krishnamurthy et al., 2024; Laskin et al., 2023; Lee et al., 2023; Shi et al., 2024; Wang et al., 2025a; Zisman et al., 2023; 2024; Song et al., 2025; Wang et al., 2025b)
-Specifically, the pretrained agent neural network takes as input not only the observation at the current time step but also an additional context.
-For example, the simplest choice of the context can be all the history of experiences that the agent has access to up to the current time step (Laskin et al., 2023).
-It is then observed that the agent’s performance improves as the length of the context grows, even if the neural network parameters are kept fixed and even if the agent is evaluated on a task that it has never encountered during the pretraining process (Laskin et al., 2023).
-This phenomenon is called ICRL.
-A widely accepted hypothesis for this performance improvement is that the forward pass of the pretrained agent network implicitly implements some RL algorithm during inference time to process the information in the context (Laskin et al., 2023; Lin et al., 2023; Wang et al., 2025a; b).
-As the context length grows, the amount of information grows, and the implemented RL algorithm in the forward pass is then able to output better actions.
-ICRL has recently received increasing attention given its cheap yet strong adaptation capability.
-The adaptation is cheap in that it does not require parameter tuning.
-So to support such adaptation, the infrastructure only needs to support the inference of the neural network.
-For example, by using ICRL in LLMs’ inference time, the LLMs are able to significantly improve on challenging tasks such as ScienceWorld (Wang et al., 2022), HMMT (Harvard–MIT Mathematics Tournament, ), and AIME (Mathematical Association of America, ) without any fine-tuning (Song et al., 2025).
-The adaptation is strong in that ICRL allows adaptation to not only unseen but also out-of-distribution tasks.
-For example, Laskin et al.
-(2023) show that after pretraining on some bandit tasks, their ICRL agent is able to adapt quickly to test bandit tasks that have opposite optimal arms to the pretraining bandit tasks.
-Kirsch et al.
-(2023) show that after pretraining on Ant robot manipulation tasks, their ICRL agent is able to adapt to pole balancing tasks.
-Despite the empirical success and the promising potential, the safety of ICRL has long been ignored.
-Here we discuss safety in the framework of constrained Markov Decision Processes (Altman, 2021), where each action the agent takes not only incurs a reward but also incurs a cost.
-Then when an agent adapts to a new task, it needs to not only maximize the rewards but also minimize the cost, all without updating the parameters.
-We argue that safe ICRL is a necessary condition to enable many possible real-world applications of ICRL, such as embodied AI.
-Although such safe decision-making problems have been well studied in the classical RL setup (Garcıa & Fernández, 2015), to our knowledge, no prior work has studied the safety of ICRL, which is the gap that this paper shall close.
-We envision that a safe ICRL agent should have the following two desiderata, both of which should be achieved without any parameter updates during the adaptation to new tasks.
-  • When evaluated in out-of-distribution tasks, the agent should be able to maximize the rewards while keeping the cost below a user-specified threshold.
-  • The agent should actively react to the cost threshold.
-For example, a smaller cost threshold should induce a more conservative behavior, while a larger cost threshold should induce a more aggressive behavior.
-The key contribution of this work is to design the first safe ICRL agent that simultaneously fulfills the two desiderata. Specifically, we make the following two contributions.
-  • We study both off-line (cf.
-(Laskin et al., 2023)) and on-line pretraining approaches (cf.
-(Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b)) for safe ICRL.
-We identify key limitations of the offline approach and propose a novel online pretraining approach that fulfills the two desiderata.
-  • We empirically validate our proposed approach in challenging safe decision making problems.
-Particularly, our test task is not only unseen during the pretraining stage but also out of the pretraining task distribution.
-As a result, to succeed in our benchmarks, the agent has to demonstrate strong out-of-distribution generalization.
-Particularly, the agent has to explore efficiently in out-of-distribution test tasks, all without parameter updates.
+(2025)) has emerged as an RL paradigm that achieves this trial-and-error process without any neural network parameter updates (Duan et al., 2016; Xu et al., 2022; Lee et al., 2023; Laskin et al., 2023; Raparthy et al., 2023; Sinii et al., 2023; Huang et al., 2024; Krishnamurthy et al., 2024).
+Specifically, ICRL enables agents to adapt to new tasks at test time by conditioning on their interaction history, without updating network parameters.
+This allows a pretrained model to generalize to unseen environments with improving performance as more contextual experience accumulates.
+Using ICRL in real-world deployments is especially desirable because, to enable adaptation, the infrastructure only needs to support the inference of the neural network.
+Despite the empirical success and the promising potential of ICRL (Laskin et al., 2023; Kirsch et al., 2023; Dai et al., 2024; Dong et al., 2024), the safety of ICRL has long been overlooked, which is a necessary condition for many possible real-world applications, such as embodied AI.
+In standard RL, agents are deployed on tasks they are trained for and expected to satisfy the safety constraints enforced during training.
+In contrast, in ICRL, agents explore and learn new tasks during test time.
+Therefore, safety constraints must be satisfied not just by the final learned behavior, but also during the adaptation process (Xu & Zhu, 2025).
+To our knowledge, no prior work has studied the safety of ICRL, which is the gap that this paper aims to close.
+We model the safe ICRL framework as a constrained Markov decision process (CMDP, Altman (2021)), where agents receive not only a reward but also a cost signal after each action.
+We envision that safe ICRL agents should maximize the rewards while keeping the cost below a user-specified cost budget even when evaluated in out-of-distribution (OOD) tasks.
+In addition, they should actively react to the cost budget; e.g., a smaller cost budget should induce more conservative behavior for safety, while a larger cost budget should allow more aggressive behavior for reward maximization.
+Meeting these desiderata is challenging; existing ICRL methods have no mechanism to incorporate cost signals, while safe meta-RL methods use past trial history only to fine-tune parameters, which does not capture the precise details of past experiences.
+The key contribution of this work is to design the first safe ICRL agent that simultaneously fulfills the aforementioned desiderata, all without any parameter updates during the adaptation to new tasks.
+Specifically, we make the following contributions:
+  • Safe ICRL Problem Formulation. We formalize safe ICRL under the CMDP framework, where agents adapt to new tasks at test time without parameter updates while adjusting to safety budgets.
+  • OOD Benchmarks for Safe In-Context Adaptation.
+We introduce benchmark environments that evaluate safe exploration and extrapolative generalization beyond interpolation regimes under safety constraints.
+  • SCARED: Safe Contextual Adaptive Reinforcement via Exact-Penalty Dual.
+We propose SCARED, an online reinforcement pretraining method that satisfies CMDP safety constraints.
+We show SCARED outperforms the baselines, increasing cumulative reward and decreasing cost as context grows, while adapting to the safety budgets.
 2 Background
-We consider a finite horizon Markov Decision Process (MDP, Bellman (1957); Puterman (2014)) with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1)$, and a horizon length $T$.
-At time step 0, an initial state $S_{0}$ is sampled from $p_{0}$.
-At time step $t$, an agent at a state $S_{t}$ selects an action $A_{t}$ according to its policy $\pi:\mathcal{A}\times\mathcal{S}\to[0,1]$, i.e., $A_{t}\sim\pi(\cdot|S_{t})$.
-The agent then proceeeds to a successor state $S_{t+1}\sim p(\cdot|S_{t},A_{t})$ and obtains a reward $R_{t+1}\doteq r(S_{t},A_{t})$.
-This process continues until the time T1T-1, where the agent executes the action $A_{T-1}$ and receives the last reward $R_{T}$.
-We use $\tau\doteq(S_{0},A_{0},R_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T})$ to denote an episode.
-We use $k$ to index episodes and $t$ to index time steps within an episode.
-When necessary, we write $S_{t}^{k}$ to denote the state at time step $t$ within the $k$-th episode.
-$A_{t}^{k}$ and $R_{t}^{k}$ are similarly defined.
-For any episode $\tau$, we use $G(\tau)\doteq\sum_{t=1}^{T}R_{t}$ to denote the return associated with this episode.
-The performance of a policy $\pi$ is then measured by the expected total rewards $J(\pi)\doteq\mathbb{E}_{\tau\sim\pi}[G(\tau)]$.
-One fundamental task in RL, called control, is to adapt the policy $\pi$ to maximize $J(\pi)$.
-Modern RL methods solve the control problem via parameterizing the policy $\pi$ with neural networks (Mnih et al., 2015).
-We denote such parameterized policy as $\pi_{\theta}$, where $\theta$ denotes the parameters of the neural network.
-Typically, $\theta$ is updated at every time step, e.g., $\theta_{t+1}$ is obtained by updating $\theta_{t}$ with the newly obtained information $(S_{t},A_{t},S_{t+1},R_{t+1})$.
-The performance of the policy $\pi_{\theta_{t}}$ improves along with such parameter updates.
-In other words, the reinforcement learning process occurs with the parameter updates.
-ICRL is an emerging paradigm that allows the reinforcement learning process to occur without any parameter updates.
-In ICRL, the policy $\pi$ takes an additional context as input.
-We denote the context at time $t$ within the $k$-th episode as $H_{t}^{k}$ and express the policy as $\pi(\cdot|S_{t}^{k},H_{t}^{k})$ to emphasize such dependency.
-The construction of the context is an active research area and we refer the reader to Moeini et al.
-(2025) for a detailed survey.
-The simplest example is to use all the history as context, e.g., $H_{t}^{k}\doteq(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},S_{1}^{k},\dots S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k})$.
-After some pretraining procedure on multiple MDPs to be detailed soon, we can obtain a parameter $\theta_{*}$.
-It is observed that the performance of $\pi_{\theta_{*}}(\cdot|S_{t}^{k},H_{t}^{k})$ improves as the context grows with $\theta_{*}$ kept fixed.
-In other words, the reinforcement learning process now occurs without any parameter updates.
-This cannot be attributed to the hypothesis that $\theta_{*}$ memorizes a good policy, as such performance improvement is also observed even when $\pi_{\theta_{*}}$ is evaluated in out-of-distribution MDPs that are very different from the pretraining MDPs (Laskin et al., 2023; Kirsch et al., 2023).
-A widely accepted hypothesis is that the neural network $\theta_{*}$ implicitly implements some RL algorithm in its forward pass, such that it can process the context with the RL algorithm in the inference time (Moeini et al., 2025).
-As the context grows, the inference-time RL algorithm in the forward pass gains access to more information.
-So the performance of the policy improves.
-This hypothesis is also theoretically backed by Lin et al.
+2.1 Markov Decision Processes
+We model the interaction with an environment for a given task as a finite-horizon MDP (Puterman, 2014).
+An MDP is defined by a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1]$, and a horizon length $T$.
+An agent interacts with the environment by following a policy $\pi:\mathcal{A}\times\mathcal{S}\to[0,1]$.
+Specifically, after starting from an initial state $S_{0}\sim p_{0}$, at any time step $t$, the agent takes an action $A_{t}\sim\pi(\cdot|S_{t})$, observes a reward $R_{t+1}\doteq r(S_{t},A_{t})$, and proceeds to a next state $S_{t+1}\sim p(\cdot|S_{t},A_{t})$.
+This process continues until the time T1T-1, when the agent takes a final action $A_{T-1}$ and receives the last reward $R_{T}$.
+We use $\tau\doteq(S_{0},A_{0},R_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T})$ to denote an episode, and use $k$ to index episodes and $t$ to index time steps in an episode.
+We write $S_{t}^{k}$, $A_{t}^{k}$, and $R_{t}^{k}$ to denote the state, action, and reward at time step $t$ in the $k$-th episode, respectively.
+For any episode $\tau$, we define its return as $G(\tau)\doteq\sum_{t=1}^{T}R_{t}$.
+The performance of a policy $\pi$ is then measured by the expected total rewards $J(\pi)\doteq\mathbb{E}_{\tau\sim\pi}[G(\tau)]$ obtained under $\pi$.
+The fundamental task in MDPs is to obtain an optimal policy $\pi$ that maximizes $J(\pi)$, which typically requires a learning-based approach when a transition model is absent.
+2.2 Reinforcement Learning
+The goal of RL is to find an optimal policy for a given task by learning through interaction with the environment (Sutton & Barto, 2018).
+Classic approaches often construct tabular value functions that are used to explicitly map each state to an action, therefore are feasible only when the state and action spaces are discrete and small.
+Recent RL techniques, combined with modern deep learning methods can handle large or continuous state–action spaces via function approximation with neural networks parameterized by weights $\theta$ (Wang et al., 2022).
+Starting from an initial policy $\pi_{\theta_{0}}$, deep RL methods typically update the parameters using observed transitions $(S_{t},A_{t},S_{t+1},R_{t+1})$ to obtain $\pi_{\theta_{t+1}}$.
+Performance can improve through these updates, which aim to reduce prediction errors via backpropagation through a loss function.
+These techniques have been successful at learning high-performing policies for a particular task; however, their performance can degrade substantially across OOD tasks and environments.
+In such settings, policies are often re-learned or fine-tuned, which limits the large-scale deployment of a single trained policy.
+2.3 In-Context Reinforcement Learning
+Recently, ICRL has emerged as a paradigm that allows RL agents to adapt to new tasks at test time without requiring any parameter updates after pretraining.
+In ICRL, the parameters $\theta_{*}$ are first learned with a pretraining procedure on multiple MDPs.
+The resulting policy $\pi_{\theta_{*}}$ parameterized by $\theta_{*}$ is then deployed to adapt to new tasks by taking context as additional input.
+We denote the context at time $t$ within the $k$-th episode as $H_{t}^{k}$ and we write the policy as $\pi(\cdot|S_{t}^{k},H_{t}^{k})$ to make this dependence explicit.
+A simple choice is to use the history of transitions as context (see Moeini et al.
+(2025) for alternative approaches); i.e., $H_{t}^{k}\doteq(\tau_{1},\dots,\tau_{k-1},\tau_{k}^{t})$ where $\tau_{k}^{t}\doteq(S_{0}^{k},A_{0}^{k},R_{1}^{k},S_{1}^{k},\dots S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k})$ is the observed prefix of the current episode.
+Adaptation to a new task in ICRL occurs as the performance of $\pi_{\theta_{*}}(\cdot|S_{t}^{k},H_{t}^{k})$ improves with growing context while $\theta_{*}$ remains fixed.
+This generalization capability is hypothesized to be a consequence of the neural network implicitly implementing an RL algorithm in its forward pass, processing the context at inference time (Laskin et al., 2023; Kirsch et al., 2023).
+As the context grows, this inference-time algorithm gains access to more information, improving policy performance.
+This hypothesis is also theoretically supported by Lin et al.
 (2023); Wang et al.
-(2025a; b).
-We now elaborate on the pretraining methods for ICRL.
-According to the taxonomy in Moeini et al.
-(2025), the pretraining can be divided into supervsied pretraining (Laskin et al., 2023; Lin et al., 2023) and reinforcement pretraining (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b; Wang et al., 2025a; b).
-Supervised pretraining is essentially behavior cloning but the goal is to imitate an algorithm instead of a policy.
-Supervised pretraining is usually done in an offline manner.
+(2025a, b).
+Despite the remarkable generalization of ICRL to OOD tasks, the safety of such generalization has been overlooked.
+We now discuss the pretraining methods for ICRL, which can typically be divided into supervised pretraining (Laskin et al., 2023; Lin et al., 2023) and reinforcement pretraining (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a, b; Wang et al., 2025a, b).
+Supervised pretraining employs behavior cloning with the goal of imitating an algorithm rather than a policy, and it is usually done in an offline manner.
 Specifically, by running an existing RL algorithm on an MDP, we can collect a sequence of episodes, denoted as $\Xi\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$.
 We call $\Xi$ a trajectory.
-Suppose the RL algorithm behaves well, we would expect that the episode return $G(\tau_{k})$ increases as $k$ increases.
-So the trajectory $\Xi$ demonstrates the learning process of the RL algorithm in the MDP.
-In supervised pretraining, we collect multiple trajectories by running multiple existing RL algorithms on multiple MDPs, yielding a dataset $\quantity{\Xi_{i}}$.
+The episode return $G(\tau_{k})$ is expected to increase as $k$ increases; therefore, the trajectory $\Xi$ can be used to demonstrate the learning progression of the RL algorithm on the MDP.
+In supervised pretraining, multiple trajectories are collected by running multiple existing RL algorithms on multiple MDPs, yielding a dataset $\quantity{\Xi_{i}}$.
 The policy $\pi_{\theta}$ is then trained with an imitation learning loss.
@@
 $-\log\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k}).$ (1)
-By asking the neural network to imitate the behavior of RL algorithms demonstrated in the dataset, we expect the neural network to be able to distill some RL algorithm into its forward pass.
-This is known as algorithm distillation (Laskin et al., 2023).
-Instead of distilling existing RL algorithms, reinforcement pretraining asks the network to discover its own RL algorithm.
+By optimizing $\theta$ to imitate the behavior of RL algorithms demonstrated in the dataset, the policy $\pi_{\theta}$ is expected to embed some RL-like mechanism into its forward pass, which is known as algorithm distillation (Laskin et al., 2023).
+Instead of distilling existing RL algorithms, reinforcement pretraining optimizes the network to maximize the return on many different MDPs, guiding it to discover its own RL algorithm by providing the interaction history.
 Reinforcement pretraining is typically done in an online manner.
@@
 The long sequence context is one of the main driving forces for the remarkable generalization capability of ICRL.
-3 Safe In-Context Reinforcement Learning
-While ICRL has demonstrated remarkable generalization in out-of-distribution tasks, the safety of such generalization has been largely overlooked.
-This paper is the first to study the safety of ICRL.
-Particularly, we use the constrained MDP (CMDP) framework (Altman, 2021).
-In addition to the reward function $r$, we also have a cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$.
-At a state $S_{t}$, the agent executes an action $A_{t}$ and receives both a reward $R_{t+1}$ and a cost $C_{t+1}\doteq c(S_{t},A_{t})$.
-An episode is then
+3 Problem Formulation: Safe ICRL
+We study the safety of ICRL using the CMDP framework (Altman, 2021).
+In addition to the reward function $r$, a CMDP involves a cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ with an associated user-given budget $\delta$.
+At each time step $t$, after taking an action $A_{t}$ at a state $S_{t}$, the agent additionally receives a cost $C_{t+1}\doteq c(S_{t},A_{t})$.
+An episode in a CMDP therefore becomes
 $\tau=(S_{0},A_{0},R_{1},C_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T},C_{T}).$ (3)
-We use $G_{c}(\tau)\doteq\sum_{t=1}^{T}C_{t}$ to denote the toal cost in the episode $\tau$.
-After some pretraining procedure, we obtain a policy $\pi_{\theta_{*}}$.
-We then execute this policy in a test MDP for multiple episodes $\tau_{1},\dots,\tau_{K}$.
-With ICRL, we can expect that the episode return $G(\tau_{k})$ increaes as $k$ increaes.
-But for safe ICRL, we would like to additionally see that the episode total cost $G_{c}(\tau_{k})$ decreases as $k$ increases, ideally below some user-given threshold, say $\delta$.
-Moreover, when the user is more tolerant of the cost (i.e., with a larger $\delta$), the agent should obtain higher rewards (i.e., larger $G(\tau_{k})$).
-We formalize these desiderata as the following constrained problem:
-$\textstyle\max_{\theta}\mathbb{E}_{\pi_{\theta}}[\sum_{k=1}^{K}G(\tau_{k})]\mbox{\quad s.t.\quad}\forall k,\mathbb{E}_{\pi_{\theta}}\quantity[G_{c}(\tau_{k})]\leq\delta.$ (4)
-Here, $K$ is the number of episodes that are allowed in the test time.
-$\tau_{k}$ is the $k$-th epsiode sampled from the policy $\pi_{\theta}$.
-The policy $\pi_{\theta}$ can depends on the context (i.e., $\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k})$) but the neural network parameters $\theta$ is fixed for all the $K$ episodes.
-We note that (4) is defined on a single test MDP that the algorithm has no access to during pretraining.
-This test MDP is usually sufficiently different from the MDPs that the algorithm has access to during pretraining.
-As a result, the agent cannot solve this test MDP via memorizing some good policies in the pretraining MDPs.
-Instead, it has to perform reinforcement learning over the $K$ episodes, although the parameters $\theta$ are fixed.
-Notably, although we formulate our problem in MDPs, the tasks we work on are only partially observable.
-This is a widely used practice in the RL community to promote the clarity of presentation (Mnih et al., 2015).
-We now explore different pretraining methods for solving the above problem.
-Safe Supervised Pretraining.
-We first establish safe supervised pretraining as a baseline approach.
-Similar to naive supervised pretraining, we collect a dataset $\quantity{\Xi_{i}}$ by executing various safe RL algorithms (e.g., Tessler et al.
-(2018); Ray et al.
-(2019)) in various safety-sensitive tasks (e.g., Ji et al.
-(2023); Gu et al.
-(2025)).
-Now each trajectory $\Xi_{i}$ consists of multiple episodes and each episode contains both the reward and the cost (cf.
-(3)).
-To further guide the action generation, we additionally condition the policy on both the return-to-go (RTG) $G_{t}(\tau)$ and the cost-to-go (CTG) $G_{c,t}(\tau)$.
-RTG is the total future rewards in the current episode, i.e., $G_{t}(\tau)\doteq\sum_{i=t+1}^{T}R_{t}$, and CTG is the total future cost in the current episode, i.e., $G_{c,t}(\tau)\doteq\sum_{i=t+1}^{T}C_{t}$.
-The RTG and CTG are inspired by Liu et al.
-(2023) for constrained decision Transformers and are designed to bias the agent towards achieving the specified return and cost.
-The main difference from Liu et al.
-(2023) is that in addition to RTG and CTG, we use entire episodes in input, while Liu et al.
-(2023) only includes states and actions but do not include per-step reward and per-step cost.
-As a result, Liu et al.
-(2023) still misses the opportunity to learn the algorithm demonstrated in the dataset, and they demonstrated only limited generalization capability.
-Specifically, given a trajectory $\Xi=(\tau_{1},\dots,\tau_{K})$ in the dataset, the safe supervised pretraining loss at a state $S_{t}^{k}$ is
-$-\log\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})).$ (5)
-We recall that the context $H_{t}^{k}$ is defined as $(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},C_{1}^{k},\dots,S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k},C_{t}^{k})$.
-In the pretraining, we do have access to RTG and CTG since the episode is already complete.
-But in the test time, we do not have access to them since they are calculated based on future rewards and costs, which are not available yet.
-Instead, we replace them with a spectrum of target RTG and CTG values.
-This gives the user the opportunity to control the trade-off between rewards and costs in the test time without fine-tuning the parameters.
-For example, in the test time, if the ratio RTG / CTG is small, it means the user is more risk-tolerant.
-If the ratio RTG / CTG is large (in the extreme case, CTG = 0), it means the user is more safety-focused.
-Although the proposed safe supervised pretraining approach is to our knowledge novel, it is indeed a straightforward extension of existing unconstrained superivsed pretraining approaches like Laskin et al.
-(2023).
-We regard this as a baseline approach for sanity check.
-Any meaningful safe ICRL method should at least outperform this baseline.
-Safe Reinforcement Pretraining.
-Our safe supervised learning follows the offline pretraining paradigm.
-It is well-known that the performance of offline trained policies is usually worse than the online trained ones when online training is plausible (Vinyals et al., 2019; Mathieu et al., 2021).
-Motivated by this, we now study safe reinforcement pretraining.
-We note that even when we consider online pretraining, we still do not have access to the test MDP where (4) is defined.
-By online pretraining, we mean that we can interact with the pretraining MDPs directly.
-By contrast, offline supervised pretraining has access to only datasets collected from pretraining MDPs.
-The underlying hypothesis of reinforcement pretraining is that by minimizing the loss (2) in a wide range of pretraining MDPs, the sequence model should be able to discover its own RL algorithm and implement it in the forward pass (see Lu et al.
-(2023); Grigsby et al.
-(2024a; b) for empirical evidence and Wang et al.
-(2025b) for theoretical evidence).
-Following this hypothesis, we now design an algorithm to solve (4) on a single pretraining MDP.
-To solve (4), primal-dual approaches in safe-RL (Achiam et al., 2017; Tessler et al., 2018; Liang et al., 2018; Ray et al., 2019) convert it to the dual problem below [^1] We omit $\theta$ and use $\max_{\pi}f(\pi)$ instead of $\max_{\theta}f(\pi_{\theta})$ for simplicity.:
-$\textstyle\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})=\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}\quantity[{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\sum_{k=1}^{K}\lambda_{k}({\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta)]$ (6)
-where ${\bm{\lambda}}\succeq 0$ is a componentwise inequality for the Lagrangian multipliers.
-Notably, the function class of the policy $\pi$ in $\max_{\pi}$ is all the policies in the form of (5) but without RTG, i.e., the policy can depend on the context and the CTG.
-However, we observe that while the CMDP formulation requires a separate constraint and Lagrangian multiplier for each evaluation episode $\tau_{k}$, this is not desirable in our problem, as it would need to fix the number of evaluation episodes at the beginning of pretraining.
-Using fewer episodes would result in some multipliers being updated fewer times than others, which makes the optimization unsable and using more episodes is not possible.
-Furthermore, there should be symmetry among all the constraints since they are all coming from a single cost function but simply correspond to different episodes.
-These encourage the use of a single multiplier for all episodes.
-However, when using a single multiplier, we would need a new way to not over-penalize the policy in the episodes that are already satisfying the constraint.
-To this end, We propose a modified Lagrangian function and an iterative optimization scheme and show it is more stable empirically while satisfying the desirable optimization properties.
-Let $g_{k}(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta$.
-We consider the following surrogate objective for the policy
-$\textstyle L_{\Sigma}(\pi,\lambda)={\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\lambda\sum_{k=1}^{K}[g_{k}(\pi)]_{+}$
+We denote the total cost of an episode $\tau$ by $G_{c}(\tau)\doteq\sum_{t=1}^{T}C_{t}$, which should, in expectation, remain below the budget $\delta$.
+The objective of ICRL is to pretrain a set of parameters $\theta_{*}$ such that the resulting policy $\pi_{\theta_{*}}$ maximizes the sum of expected returns for a new task at test time.
+This is usually achieved by the return $G(\tau_{k})$ of an episode $\tau_{k}$ being increased with $k$, as the policy learns more about the environment.
+In safe ICRL, we additionally expect the episode cost $G_{c}(\tau_{k})$ to decrease as $k$ increases, ideally remaining below a user-specified cost budget $\delta$.
+We formalize this safe ICRL objective as follows:
+$\textstyle\max_{\theta}\mathbb{E}_{\pi_{\theta}}[\sum_{k=1}^{K}G(\tau_{k})]\text{~~s.t.~~}\forall k,\mathbb{E}_{\pi_{\theta}}\quantity[G_{c}(\tau_{k})]\leq\delta.$ (4)
+Here, $K$ is the number of total test episodes, and $\tau_{k}$ denotes the $k$-th test episode obtained by executing the policy $\pi_{\theta}$ on a novel CMDP that is not accessible during pretraining.
+Throughout test time, the CMDP and the policy parameters $\theta$ are fixed; therefore, the policy $\pi_{\theta}(\cdot|S_{t}^{k},H_{t}^{k})$ can adapt to the CMDP only by utilizing the context, i.e., the history of test transitions.
+Although we formulate our problem using CMDPs, our formulation can be used for partially observable systems without loss of generality.
+4 SCARED: Safe Contextual Adaptive Reinforcement Pretraining
+We now introduce SCARED, our online reinforcement pretraining approach for the safe ICRL problem.
+SCARED enables safe in-context adaptation to unseen CMDPs by regulating episode-level costs within a given budget.
+We establish a pretraining procedure that minimizes the loss (2) across a range of CMDPs, extending the generalization capability from reward maximization to also respecting the safety budget, solely based on context without parameter updates at test time.
+Specifically, we design an algorithm to solve (4) by converting the primal problem to its dual form [^1] We omit $\theta$ and write $\max_{\pi}f(\pi)$ instead of $\max_{\theta}f(\pi_{\theta})$ for brevity.:
+$\hskip-5.0pt\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})=$
+$\hskip-5.0pt\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}\quantity[{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}{-}\sum_{k=1}^{K}\hskip-1.0pt\lambda_{k}({\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}{-}\delta)].$ (5)
+where ${\bm{\lambda}}\succeq 0$ denotes componentwise nonnegativity of the Lagrangian multipliers.
+The maximization over $\pi$ ranges over contextual policies that condition not only on the interaction history but also on the total cost expected over the remainder of the episode.
+We refer to this total future cost as cost-to-go (CTG) and define it as $G_{c,t}(\tau)\doteq\sum_{i=t+1}^{T}C_{i}$.
+At the start of each episode, CTG is set to the budget $\delta$, i.e., $G_{c,0}=\delta$.
+Policies then sample actions conditioned on CTG as $A_{t}^{k}\sim\pi_{\theta}\!\left(\cdot\mid S_{t}^{k},H_{t}^{k},G_{c,t}(\tau_{k})\right)$.
+In safe ICRL, applying a separate constraint and Lagrange multiplier to each evaluation episode $\tau_{k}$ is undesirable as this requires fixing the number of episodes at the start of pretraining and leads to uneven update frequencies across multipliers.
+For example, if we expect $K$ evaluation episodes, we must initialize at least $K$ multipliers.
+If pretraining rollouts contain fewer episodes than the number of multipliers, later multipliers are updated less frequently, which destabilizes the optimization (see Figure 5.(c)).
+Furthermore, the constraints should be symmetric since they are associated with a single cost function but simply correspond to different episodes.
+These motivate the use of a single multiplier for all episodes.
+However, using a single multiplier will over-penalize the policy on episodes that already satisfy the constraint.
+To prevent this, we propose a modified Lagrangian function and an iterative optimization scheme.
+We show this approach is empirically more stable while preserving the desirable optimization properties.
+Let $g_{k}(\pi)\coloneqq{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}-\delta$. We consider the following surrogate objective for the policy
+$\textstyle L_{\Sigma}(\pi,\lambda)={\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}-\lambda\sum_{k=1}^{K}[g_{k}(\pi)]_{+}$
 where $x]_{+}\doteq\max\quantity{x,0}$ and perform the following iterative updates:
-$\textstyle\pi_{t+1}\in\operatorname*{argmax}_{\pi}L_{\Sigma}(\pi,\lambda_{t})\,\quad\lambda_{t+1}=[\lambda_{t}+\eta\,\max_{k}g_{k}(\pi_{t+1})]_{+}\quad\eta>0.$ (7)
+$\textstyle\pi_{t+1}\in\operatorname*{argmax}_{\pi}L_{\Sigma}(\pi,\lambda_{t})$ (6)
+$\lambda_{t+1}=[\lambda_{t}+\eta\,\max_{k}g_{k}(\pi_{t+1})]_{+}\quad\eta>0.$
+We analyze the updates in (6) under the following regularity assumptions.
 Assumption 1.
-The expected return $\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]$} and expected cost $\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]$} are bounded and continuous in $\pi$.
+The expected return $\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]$} and expected cost $\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]$} are bounded and continuous in $\pi$.
 The constrained problem (4) admits an optimal feasible policy $\pi^{\star}$.
 Assumption 2.
-[Paternain (2018)] Problem (4) satisfies conditions ensuring zero duality gap with (6) and there exists optimal Lagrange multipliers ${\bm{\lambda}}^{\star}\succeq 0$ such that
-$\textstyle\underset{\text{s.t.}\ \forall k,{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}\leq\delta}{\max_{\pi}}{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}\ \;=\;\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})$
-We now show that the set of the fixed points of (7) and the set of the optimizers of (4) are equal.
+[Paternain (2018)] Problem (4) satisfies conditions ensuring zero duality gap with (4) and there exist optimal Lagrange multipliers ${\bm{\lambda}}^{\star}\succeq 0$ such that
+$\underset{\text{s.t.}\ \forall k,{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}\leq\delta}{\max_{\pi}}{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}\ =\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})$
+The following theorem shows that the set of the fixed points of (6) and the set of the optimizers of (4) are equal, meaning that the objective of our approach aligns with the safe ICRL problem statement.
 Theorem 1.
@@
 Let Assumptions 1 - 2 hold.
-Then every fixed point $(\bar{\pi},\bar{\lambda})$ of (7) is feasible and primal-optimal for (4).
-Conversely, every primal-optimal policy $\pi^{*}$ admits a corresponding fixed point $(\pi^{*},\bar{\lambda})$ for any multiplier $\bar{\lambda}\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$, where $\bm{\lambda}$$^{\star}$ is an optimal dual solution.
-To implement the policy update (7), we consider an actor-critic approach following Grigsby et al.
-(2024a).
-The pseudocode is provided in Algorithm 1.
-Since our update (7) resembles an exact penalty convex optimization technique when $\lambda\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$ (Nocedal & Wright, 2006), we call the resulting actor-critic algorithm Exact Penalty Policy Optimization (EPPO).
+Then every primal-optimal policy $\pi^{*}$ admits a corresponding fixed point $(\pi^{*},\bar{\lambda})$ for any multiplier $\bar{\lambda}\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$, where $\bm{\lambda}$$^{\star}$ is an optimal dual solution.
+Conversely, every fixed point $(\bar{\pi},\bar{\lambda})$ of (6) is feasible and primal-optimal for (4).
+Our update (6) resembles an exact penalty convex optimization technique when $\lambda\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$ (Nocedal & Wright, 2006), hence the naming.
+We implement it with an actor-critic approach following Grigsby et al.
+(2024a); see Algorithm 1 for pseudocode.
+Separate critics estimate the reward Q-function $Q_{\theta_{v}}$ and cost Q-function $Q^{c}_{\theta_{c}}$, trained with TD targets from target networks.
+The actor maximizes $Q_{\theta_{v}}$ while penalizing $Q^{c}_{\theta_{c}}$ for episodes exceeding the budget.
 5 Experiments
-We now empirically investigate the proposed safe ICRL algorithms. Specifically, we aim to answer the following two questions.
+We now empirically investigate the proposed safe ICRL algorithms, aiming to answer the following research questions:
   • (i)
-    Do the proposed safe ICRL algorithms demonstrate generalization to out-of-distribution (OOD) tasks in complex environments with safety constraints?
+    Does SCARED generalize to both out-of-distribution tasks and unseen in-distribution (ID) tasks in safety-constrained environments?
   • (ii)
-    How effectively do the proposed safe in-context RL algorithms achieve flexible reward-cost tradeoffs under varying cost tolerance?
-We evaluate our approach on two constrained environments: (1) SafeDarkRoom and (2) SafeDarkMujoco: Point and Car.
-These are modified from DarkRoom (Laskin et al., 2023) and SafetyGym benchmark (Ji et al., 2023) respectively.
+    How effectively does SCARED achieve flexible reward-cost tradeoffs under varying safety budgets at test time, without parameter updates?
+5.1 Baselines
+We consider two baseline classes: algorithm distillation for safe RL, which encodes safety into the policy’s forward pass without test-time parameter updates, and safe meta-RL methods, which adapt via parameter updates at evaluation but do not use in-context interaction histories for action selection.
+Algorithm Distillation for Safe RL.
+We adopt algorithm distillation for safe RL (safe AD) as a strong baseline for safe ICRL.
+Following algorithm distillation (Laskin et al., 2023), we collect a dataset $\{\Xi_{i}\}$ by executing existing safe RL algorithms (e.g., (Tessler et al., 2019; Ray et al., 2019)) across safety-constrained environments (e.g., (Ji et al., 2023; Gu et al., 2025)).
+As a result, each trajectory $\Xi_{i}$ consists of multiple episodes and each episode contains both the reward and the cost (cf.
+(3)).
+To ensure a fair and competitive baseline, the policy is conditioned on the full in-context interaction history $H^{k}_{t}$, which includes per-step reward and cost signals, as well as on CTG and return-to-go (RTG), a target specifying the desired future reward that steers the policy toward a desired reward level (Liu et al., 2023).
+The raw reward and cost information in $H^{k}_{t}$ enables algorithm distillation (Laskin et al., 2023), allowing the model to imitate adaptation behavior present in the offline trajectories.
+In contrast, RTG and CTG serve a distinct role as conditioning variables that bias action selection toward desired reward-cost tradeoffs at test time.
+Formally, given a trajectory $\Xi=(\tau_{1},\dots,\tau_{K})$, the safe AD objective at state $S_{t}^{k}$ is $\log\pi_{\theta}\!\left(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)$.
+During pretraining, RTG and CTG are available since episodes are complete.
+At test time, they are replaced by user-specified targets, allowing control over the reward-cost tradeoffs without parameter updates.
+Safe AD relies on a large replay buffer of logged learning histories, which must be subsampled due to context-length constraints and labeled with RTG to distinguish between early, intermediate, and expert phases of training (Dai et al., 2024).
+We also evaluate a noise-based variant of Safe AD (Zisman et al., 2023); however, training on optimal trajectories with noise fails to provide the behavioral diversity required for OOD adaptation, highlighting the learning mechanism in our Safe AD baseline for reliable in-context safety adaptation (Appendix D).
+Safe Meta RL.
+We compare our approach against gradient-based safe meta RL methods, specifically MAML with penalty (Finn et al., 2017), and SafeMeta (Xu & Zhu, 2025).
+MAML with penalty optimizes a model-agnostic initialization via cost-weighted penalties that can be adapted to new tasks with a few gradient steps.
+SafeMeta combines a closed-form, one-step safe policy adaptation method with a Hessian-free meta-training algorithm to guarantee anytime safety.
+By design, these methods adapt through per-task parameter updates rather than conditioning on cross-episode interaction histories.
+5.2 Experimental Setup
+Safe ICRL Benchmarks.
+We evaluate our approach on three constrained environments, categorized into two complementary generalization regimes: (i) structural OOD tasks and (ii) unseen ID tasks.
+First, we use SafeDarkRoom and SafeDarkMujoco (Point and Car) to evaluate adaptation to OOD tasks with shifted goal and obstacle configurations.
+Second, we evaluate on SafeVelocity (HalfCheetah and Ant), representing the unseen ID task regime commonly used in meta-RL benchmarks.
+We discuss the difficulty gap between these two categories in Section 5.3.
+The first two environments are modified from DarkRoom (Laskin et al., 2023) and SafetyGym benchmark (Ji et al., 2023) respectively.
 SafeDarkRoom is a grid-world setting where the agent can only perceive its own position and lacks visibility of the goal or obstacle locations.
-Rewards are sparse, with a positive reward obtained upon reaching the single goal in the map.
+Rewards are sparse, with a positive reward obtained upon reaching the goal in the map.
 Costs arise from multiple obstacles scattered across the environment.
@@
 To succeed, the agent must learn obstacle positions based on encountered cost signals, introducing additional complexity compared to goal-only environments due to the presence of multiple hazards.
-SafeDarkMujoco environment operates in continuous space with MuJoCo simulation (Todorov et al., 2012) but with a setup analogous to SafeDarkRoom.
-The robot senses its internal physical states, such as velocity, acceleration, position, and rotation angle, while all lidar inputs are turned off, preventing detection of obstacles and the goal.
-Consequently, the agent must develop exploratory behaviors to identify obstacles and goals, akin to SafeDarkRoom design, while navigating to the goal and minimizing collisions by learning from the previous collisions in the context.
-Measuring OOD generalization.
+SafeDarkMujoco operates in continuous space with MuJoCo simulation (Todorov et al., 2012), with a setup analogous to SafeDarkRoom.
+The robot senses its internal physical states, such as velocity, acceleration, position, and rotation angle, while range-based sensing (e.g., lidar) is unavailable, preventing direct detection of obstacles and the goal.
+Consequently, the agent must develop exploratory behaviors to identify obstacles and goals, akin to the SafeDarkRoom design, while navigating to the goal and minimizing collisions by learning from previous collisions in the context.
+SafeVelocity follows the design of Ji et al.
+(2023); however, we vary the target velocities and hold out a set of unseen velocities for evaluation.
+Such velocity-conditioned setups are often considered within meta-RL frameworks (Finn et al., 2017), where different target velocities are sampled from a fixed range (e.g., $[0,1]$).
+We thus consider SafeVelocity tasks as unseen ID tasks, which are more straightforward than OOD generalization.
+Training Setup.
+For OOD task environments, all algorithms are trained on source tasks with center-oriented obstacles and goals.
+For unseen ID task environments, we hold out a set of target velocities multipliers uniformly sampled from $[0.5,1.0]$ in increments of $0.05$ during training and evaluating performance on these held-out velocities at test time.
+Safe AD is trained on trajectories generated using PPO-Lagrangian (Ray et al., 2019).
+For SCARED, we perform online reinforcement pretraining for 30,000 steps on SafeDarkRoom and 10,000 steps on SafeDarkMujoco using Algorithm 1, and we train for 1,000 steps on the SafeVelocity environments (HalfCheetah and Ant).
+Safe meta-RL baselines are trained for up to 15,000 steps on each environment, with most converging earlier, using their respective update rules.
+For safe AD and safe meta-RL baselines, we perform hyperparameter optimization to ensure fair comparison across methods.
+Additional implementation details and used hyperparameters are provided in Appendix C.
+Evaluation Setup.
+To answer the aforementioned research questions, we use a consistent setup for CTG and RTG during evaluation.
+For SCARED, we evaluate robustness to safety budgets by conditioning on CTG values sampled uniformly from $[1,10]$ in SafeDarkRoom, $[10,50]$ in SafeDarkMujoco, and $[0,5]$ in SafeVelocity, reporting performance averaged over 100 distinct CTG targets.
+This validates that pretraining with diverse CTG targets enables budget-conditioned generalization to OOD tasks.
+The same intervals and sample counts are used for the cost budgets of SafeMeta and MAML with penalty.
+For Safe AD, we condition the policy on paired target values.
+Specifically, we initialize RTG at $0.1$ and CTG at $0.0$, and then adjust RTG according to $\mathrm{RTG}=\max\left\{0.1,\frac{\mathrm{CTG}}{\mathrm{CTG}_{\max}}\right\}$ as CTG increases from 0 to $\mathrm{CTG}_{\max}$, where $\mathrm{CTG}_{\max}=10$ for SafeDarkRoom, $50$ for SafeDarkMujoco, and $5$ for SafetyVelocity.
+Performance is reported as an average over 100 paired RTG-CTG targets.
+5.3 Measuring OOD generalization
 Goal discovery-oriented environments such as DarkRoom are widely used in previous ICRL works (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025).
@@
 The agent spawns at the map center, with obstacles and the goal distributed proportionally closer to this center during pretraining.
-During evaluation, obstacles and the goal follow an edge-oriented distribution.
-This approach applies similarly to the SafeDarkRoom and SafeNavigation environments.
+During evaluation, obstacles and goals follow an edge-oriented distribution.
+This approach applies similarly to the SafeDarkRoom and SafeDarkMujoco environments.
 We argue that this setup is more challenging than those in prior studies (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025), as the agent must learn extrapolation rather than interpolation.
@@
 Similarly, the KL divergence holds $\lim_{\alpha\rightarrow\infty}D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)=\infty$.
+Proposition 1 shows that as $\alpha\to\infty$, the overlap between the training and test distributions vanishes.
+As a result, the model must generalize to inputs far outside the regions observed during training.
+5.4 Analysis
 Question (i).
-We now demonstrate emergent safe learning behaviors in OOD test tasks, thus giving an affirmative answer to Question (i).
-We use supervised pretraining and reinforcement pretraining in Section 3 on pretraining tasks with center-oriented obstacles and goals.
-Particularly, for supervised pretraining, the source algorithm to generate the dataset is PPO-Lagragian (Ray et al., 2019).
-goals.
-For reinforcement pretraining, we update the model for 30,000 steps on SafeDarkRoom and 10,000 steps on SafeDarkMujoco using Algorithm 1.
-After pretraining, we evaluate the learned agents in evaluation tasks with edge-oriented obstacles and goals.
-The agent is able to interact with the evaluation task for multiple episodes, but cannot update its network parameters.
-During evaluation, for supervised pretraining, we set RTG to 1.0 and CTG to 0.0, targeting successful performance while minimizing cost violations.
-Reinforcement pretraining is solely conditioned on CTG.
-To evaluate robustness to CTG values during evaluation, we set CTG to a value uniformly sampled from the interval $[1,15]$ for SafeDarkRoom and $[10,50]$ for SafeDarkMujoco.
-Precisely, we sample $100$ distinct CTG values and report the average performance over these $100$ distinct CTG values.
-We recall that although the safe supervised pretraining method we propose is novel, we mostly regard it as a straightforward baseline.
-As can be seen in Figure 1, safe supervised pretraining succeeds in SafeDarkRoom as the episode return grows and the episode cost decreases.
-This aligns with the previously observed success of algorithm distillation in DarkRoom (Laskin et al., 2023; Zisman et al., 2023).
-But this baseline approach fails in SafeDarkMujoco.
-We conjecture that this is because of insufficiency in the dataset, as supported by ablation studies (Figure 6).
-By contrast, our safe reinforcement pretraining approach consistently succeeds in all three tested domains.
-It is also worth highlighting that in SafeDarkRoom, the safe reinforcement pretraining enables much faster adaptation to the new evaluation task than the safe supervised pretraining approach.
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
-Figure 1: Evaluation performance of superivsed and reinforcement pretraining.
-The curves are averaged over 1616 distinct OOD evaluation tasks, with each task featuring edge-oriented goal and obstacles.
-The shaded regions indicate standard errors.
-The $x$-axis is the episode index $k$ during the evaluation task.
-The $y$-axis is the corresponding episode return $G(\tau_{k})$ and the epsiode cost $G_{c}(\tau_{k})$ respectively.
-The straightforward safe supervised pretraining baseline succeeds only in SafeDarkRoom, while our novel safe reinforcement pretraining method succeeds in all three domains.
+Figure 1: Evaluation performance under increasing in-context episodes.
+Curves are averaged over environment-specific OOD evaluation tasks; shaded regions denote standard errors.
+The $x$-axis shows the episode index $k$, and the $y$-axis reports the episode return $G(\tau_{k})$ (top) and episode cost $G_{c}(\tau_{k})$ (bottom).
+SCARED shows consistent improvement in return and reduction in cost as the in-context interaction grows across all environments, and outperforms all baselines.
+Safe AD shows environment-dependent performance, failing in the SafeVelocity domain and performing competitively in others, but does not achieve the same performance as SCARED.
+Safe meta-RL algorithms fail to achieve safe learning behavior due to a lack of in-context interaction, despite updating parameters during evaluation.
+We now demonstrate emergent safe learning behaviors in the test environments, thus giving an affirmative answer to Question (i).
+For OOD evaluation, we use edge-oriented obstacles and goals in SafeDarkRoom and SafeDarkMujoco.
+For unseen ID tasks, we evaluate on SafeVelocity (HalfCheetah and Ant), where target velocities are unseen during training.
+During evaluation, agents interact with each task for multiple episodes without updating parameters (except for meta-RL baselines, which adapt via parameter updates).
+Figure 1 shows SCARED consistently outperforms all baselines by learning safe behavior through in-context interaction, showing increasing episode return and decreasing episode cost as the interaction history grows across all evaluated environments.
+Safe AD performs competitively in SafeDarkRoom, showing a similar learning pattern but with slower and weaker adaptation than SCARED.
+However, its performance degrades as environment complexity increases.
+In SafetyAnt, episode cost rises while return declines through in-context episodes, and in SafeDarkMujoco, although return improves, cost does not decrease, failing safe adaptation.
+Among safe RL baselines, MAML with penalty fails to achieve safe and high-performing adaptation across the evaluated environments.
+In contrast, SafeMeta achieves higher returns in SafetyVelocity domains but does not reduce safety violations, even when allowed to update parameters during evaluation.
+This suggests that parameter adaptation alone, without cross-episode interaction histories, is insufficient to regulate safety during online exploration.
 Question (ii).
-We now demonstrate that we can adjust the behavior of the pretrained policy in test time by simply changing the RTG and/or CTG without making any parameter updates.
-Intuitively, with a higher CTG (i.e., a higher cost tolerance), the agent should afford bolder exploration and thus obtain higher rewards.
-Conversely, a lower CTG should promote conservative policies that prioritize safety, potentially at the expense of suboptimal rewards.
-For the supervised pretrained model, we initialize RTG at 0.5 and CTG at 0.0, then adjust RTG = $\max\{0.5,\frac{\text{CTG}}{10}\}$ as CTG increases from 5 to 15 for SafeDarkRoom and 0 to 10 for SafeDarkMujoco.
-This setup enables gradual increases in both RTG and CTG to demonstrate balanced reward-cost trade-offs, consistently applied to SafeDarkRoom and SafeDarkMujoco.
-However, the supervised pretrained model fails to learn trade-off behaviors, exhibiting random patterns with varying RTG and CTG (Figure 2.(b)).
-We hypothesize that this stems from insufficient CTG diversity during training, leading to dataset inadequacy.
-In our reinforcement pretrained model, we focus on controlling the CTG exclusively.
-This design simplifies the regulation of RTG based on the CTG.
-Simultaneously controlling both RTG and CTG is challenging, as determining the feasibility of RTG/CTG pairs is complex (Liu et al., 2023).
-Instead, the reinforcement pretrained model is designed to control CTG and pursue the maximal possible return for a given CTG, eliminating the need to justify the feasibility of specific solutions.
-Figure 2.(a) illustrates this relationship, showing that as cost limits increase, the return also increases.
-Supervised pretraining depends on a long replay buffer to learn from logged learning histories.
-Due to context length constraints, we need to subsample trajectories from these learning histories for supervised pretraining.
-These histories are sampled from distinct learning phases: early (low return, high cost), middle (average return, average cost), and expert (high return, low cost).
-To make sequence modeling feasible, trajectories need to be labeled with RTG, which provides information about the expected phase (Dai et al., 2024).
-However, this approach requires significantly more data, as noted by Liu et al.
-(2023).
-In contrast, reinforcement pretraining leverages online interaction histories, eliminating the need to distinguish between different data phases.
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
-Figure 2: Evaluation performance of supervised and reinforcement pretraining with varying CTG.
-The curves are displayed across a range of cost limits.
-For each cost limit, the result is averaged over 1616 distinct OOD evaluation tasks, with each task incorporating edge-oriented goals and obstacles.
-The shaded regions indicate standard errors.
-The $x$-axis is the CTG.
-The $y$-axis is the total episode return (i.e., $\sum_{k=1}^{K}G(\tau_{k})$) and the maximum episode cost (i.e., $\max_{k\in K}G_{c}(\tau_{k})$) with the corresponding CTG.
-The policy from safe reinforcement pretraining succeeds in converting a higher cost limit (i.e., a higher CTG) to a higher return while the policy from safe supervised pretraining fails to do so.
+Figure 2: Evaluation performance under varying CTG targets.
+Results are averaged over environment-specific OOD evaluation tasks; shaded regions denote standard errors.
+The $x$-axis shows the target CTG, and the $y$-axis reports the total return $\sum_{k=1}^{K}G(\tau_{k})$ (top) and the maximum episode cost $\max_{k\in[1,K]}G_{c}(\tau_{k})$ (bottom).
+SCARED shows a clear pattern: as the cost budget increases, it achieves higher total return across all environments.
+Other algorithms fail to present this budget-adjustment behavior.
+We demonstrate that the behavior of the pretrained policy can be adjusted at test time by varying CTG, without updating parameters.
+Intuitively, higher CTG values correspond to greater cost tolerance and enable more aggressive exploration and higher return.
+Conversely, lower CTG values induce more conservative, safety-focused behavior, potentially at the expense of suboptimal rewards.
+To study this, we use the same range of CTG introduced before.
+SCARED controls CTG exclusively and pursues the maximal achievable return subject to the specified cost budget, avoiding the need to reason about the feasibility of arbitrary RTG-CTG pairs, which is known to be challenging in constrained settings (Liu et al., 2023).
+Figure 2 shows that increasing the cost budget consistently leads to higher return, indicating that SCARED learns to map the cost budget to an exploration strategy during in-context adaptation.
+Notably, maximum episode costs track the CTG target closely, with nearly all episodes respecting the specified budget.
+In contrast, Safe AD fails to adapt consistently to varying CTG values, showing inconsistent behavior.
+For example, in SafeDarkRoom, Safe AD achieves relatively high total return at low CTG targets but shows a downward trend as the cost budget increases, failing to control return-cost tradeoffs.
+Similarly, safe meta-RL baselines do not show comparable budget-adjustment behavior.
+Regardless of specified CTG targets, their performance does not vary with the cost budgets, suggesting that parameter adaptation alone does not provide test-time control over reward-cost tradeoffs.
 Ablations.
-We further conduct ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, alongside specific factors such as the dataset size for supervised pretraining.
-Our findings show distinct sensitivities to these factors.
-Reinforcement pretraining is largely unaffected by model size, whereas model size significantly influences supervised pretraining performance.
-For context length, reinforcement pretraining performs better with longer sequences, while supervised pretraining performs worse.
-Conversely, supervised pretraining excels with shorter context lengths, while reinforcement pretraining performs poorly.
-These results suggest that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, handles long-term credit assignment more effectively.
-We confirm the well-established finding that supervised pretraining is highly sensitive to dataset size.
-When the dataset is small, the model fails to learn, exhibiting random behavior on out-of-distribution data.
-In Figure 5.(c), we compare EPPO to a naive primal–dual solver for (6) with per-episode multipliers $\lambda$.
-The overall structure of this naive algorithm resembles RCPO (Tessler et al., 2018).
-EPPO trains more stably with minimal tuning and solves SafeDarkRoom and SafeDarkMujoco without changes beyond time-limit related settings (Table 1).
-The detailed evaluation curves of the ablation study are provided in Figures 56 in Appendix E.
-We evaluate a variant of safe supervised pretraining incorporating algorithm distillation with noise in Appendix D.
-The results indicate that learning histories generated by noisy actions do not effectively represent true learning histories in complex environments with safety constraints.
+We conduct ablations on SCARED and Safe AD over shared factors (context length, model size) and method-specific factors (safe AD dataset size), with results reported in Appendix E, revealing distinct sensitivities to these factors.
+SCARED is largely unaffected by model size, whereas model size significantly influences safe AD performance.
+For context length, SCARED performs better with longer sequences, while safe AD performs worse.
+Conversely, safe AD excels with shorter context lengths, while SCARED performs poorly.
+These results suggest that long-term credit assignment is harder in offline RL due to dataset constraints, while online learning benefits from direct environment interaction.
+We confirm that Safe AD is highly sensitive to dataset size: with limited data, it fails to learn and shows random behavior on OOD tasks.
 6 Related Works
 ICRL.
-Learning to improve on a new MDP by interacting with it and without parameter updates is first studied in meta RL (Duan et al., 2016; Wang et al., 2016).
+Learning to improve on a new MDP by interacting with it without parameter updates was first studied in meta-RL (Duan et al., 2016; Wang et al., 2016).
 See Beck et al.
-(2025) for a detailed survey of meta RL.
-In general, the demonstrated zero parameter update generalization is weak in meta RL literature.
-Most of them rely on task identification, i.e., identifying pretraining tasks that are similar to the evaluation task and acting as if the evaluation task were the pretraining tasks.
+(2025) for a detailed survey of meta-RL.
+In general, zero-parameter-update generalization in meta-RL is limited, as most methods rely on task identification—matching evaluation tasks to similar pretraining tasks, i.e., identifying pretraining tasks that are similar to the evaluation task and acting as if the evaluation task were the pretraining tasks.
 Laskin et al.
-(2023) coined the word ICRL and demonstrated generalization to evaluation tasks that are far away from the pretraining tasks, which spurs increasing attention in ICRL.
+(2023) coined the term ICRL and demonstrated strong generalization to tasks far from the pretraining distribution, spurring growing interest in the area.
 Variations of algorithm distillation have been proposed for efficient in-context reinforcement learning, including the Decision Pretrained Transformer (Lee et al., 2023) and Algorithm Distillation with Noise (Zisman et al., 2023).
 Both approaches aim to efficiently train transformer models for in-context learning using optimal policies.
+LLMs have been shown to exhibit ICRL as an emergent capability, enabling them to keep improving on tasks when previous attempts are given along with a reward in the context (Song et al., 2025).
 See Moeini et al.
 (2025) for a detailed survey of ICRL.
-The main novelty of this work is to first study safe ICRL within the constrained MDP framework.
+However, none of these works have investigated safety within the context of ICRL.
 Safe RL.
-Safe RL methods commonly adopt CMDP formulations to enforce compliance with safety constraints during exploration and policy optimization (Garcıa & Fernández, 2015; Gu et al., 2022; Wachi et al., 2024).
-Constrained Policy Optimization (CPO) serves as a foundational algorithm in this area, effectively balancing performance rewards with safety requirements (Achiam et al., 2017; Wachi & Sui, 2020).
-Subsequent extensions, such as Constrained Update Projection and Constrained Reward Policy Optimization, have built upon CPO to enhance theoretical guarantees and practical applicability in safe RL scenarios (Xu et al., 2021; Yang et al., 2022).
+Safe RL methods commonly adopt CMDP formulations to enforce compliance with safety constraints during exploration and policy optimization (García & Fernández, 2015; Gu et al., 2022; Wachi et al., 2024).
+Constrained Policy Optimization (CPO) serves as a foundational algorithm in this area, balancing rewards with safety (Achiam et al., 2017; Wachi & Sui, 2020).
+Building on CPO, subsequent works enhance theoretical guarantees and scalability through primal-dual formulations and efficient projection-based updates (Tessler et al., 2019; Xu et al., 2021; Yang et al., 2022).
 Shielding with function encoders and conformal prediction has been proposed to handle unseen OOD environments (Kwon et al., 2025).
-However, this work emphasizes runtime adaptation rather than learning algorithms through contextual information.
-Reward-Constrained Policy Optimization (Tessler et al., 2018) is the closest work to EPPO, but it enforces safety only in expectation rather than at the episode level.
-From the offline RL perspective, the Constrained Decision Transformer (Liu et al., 2023) leverages the transformer architecture for safe RL.
+However, this work emphasizes runtime adaptation rather than learning algorithms through interaction.
+From the offline RL perspective, Constrained Decision Transformer (Liu et al., 2023) uses Transformer architecture for safe RL.
 However, existing approaches enforce safety constraints only when the evaluation task closely matches the pretraining tasks.
-In contrast, our safe ICRL framework verifies safety constraints even on evaluation tasks that differ substantially from the pretraining distribution.
+In contrast, our safe ICRL framework satisfies safety constraints even on evaluation tasks that differ substantially from the pretraining distribution.
 Safe Meta RL.
-Safe meta RL has emerged to enable fast adaptation to new tasks while enforcing safety constraints.
-Earlier work (Luo et al., 2021) uses a three-phase approach that meta-learns a safety critic from offline data across multiple environments, adapts it to a new environment with a smaller offline dataset, and employs the adapted safety critic with a recovery policy to ensure safe learning while minimizing constraint violations.
+Safe meta RL enables fast adaptation to new tasks, enforcing safety constraints.
+For example, Luo et al.
+(2021) propose a three-phase strategy that meta-learns a safety critic across environments, adapts it to new tasks using limited offline data, and use it with a recovery policy to reduce constraint violations during learning.
 Khattar et al.
@@
 (2024) proposes a cost-aware context encoder that uses supervised cost relabeling and contrastive learning to infer tasks based on safety constraints.
-To our knowledge, all previous safe meta RL works require parameter updates to achieve safe adaptation in evaluation tasks.
-By contrast, our work on safe ICRL enables safe adaptation to evaluation tasks without parameter updates.
+Xu & Zhu (2025) proposes a safe meta-RL algorithm that integrates a closed-form, one-step safe policy adaptation mechanism to provide anytime safety guarantees.
+Unlike safe meta-RL methods, which rely on parameter updates to achieve safe adaptation during evaluation, safe ICRL enables adaptation without test-time parameter updates.
 7 Conclusion
-This work pioneers the study of safe ICRL, examining both reinforcement pretraining and supervised pretraining.
-We develope a safe supervised pretraining appraoch based on the constrained decision transformer, showcasing its capability for OOD generalization while revealing its dependence on dataset size.
-We introduce EPPO, a safe reinforcement pretraining approach via exact penalty policy optimization to enforce strict cost constraints within each episode.
-Theoretically, we prove that EPPO’s fixed point satisfies strict safety constraints while maximizing rewards.
-We conduct a thorough empirical study of the proposed approaches in challenging benchmarks that provably demand OOD generalization (Proposition 1).
-Such demand for OOD generalization is missing in many benchmarks used by previous works (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025) Through those challenging benchmarks, we confirm that our pretrained safe ICRL agents adapt to not only unseen but also OOD evaluation tasks while respecting the safety constraints.
-We envision this work guiding the development of robust, generalizable, and safe RL algorithms for real-world applications.
+This work pioneers the study of safe ICRL, where agents must adapt to new and potentially OOD tasks without parameter updates while adhering to safety constraints.
+We show that existing approaches, including algorithm distillation for safe RL, and safe meta-RL baselines, are limited in this setting, as they lack mechanisms for regulating safety through in-context interaction histories and for adjusting to cost budgets during adaptation.
+To address these limitations, we introduced SCARED, a reinforcement pretraining approach for safe in-context adaptation under the CMDP framework.
+SCARED enables agents to balance reward maximization and cost minimization during adaptation by regulating episode-level safety constraints and adjusting behavior in response to different cost budgets.
+Theoretically, we prove that fixed points of SCARED optimization satisfy strict safety constraints while maximizing rewards.
+To rigorously test OOD generalization, we introduce challenging benchmarks for safe ICRL (Proposition 1), addressing a gap in prior evaluations (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025).
+Through these challenging benchmarks, we confirm that our pretrained safe ICRL agents adapt to not only unseen ID but also OOD evaluation tasks while respecting safety constraints.
+We envision this work guiding the development of robust, generalizable, and safety-aware in-context learning methods for real-world applications.
 Acknowledgments
 This work is supported in part by the US National Science Foundation under the awards III-2128019, SLES-2331904, CAREER-2442098, and CCF-1942836, the Commonwealth Cyber Initiative’s Central Virginia Node under the award VV-1Q26-001, and an Nvidia academic grant program award.
+Broader Impacts Statement
+This work introduces the first study of safety in ICRL, a paradigm in which agents adapt to new tasks at deployment time without parameter updates.
+By providing a framework for enforcing user-specified safety budgets during test-time adaptation, this work is relevant to real-world applications such as robotics and embodied AI, where retraining is impractical and unsafe exploration can have real consequences.
+The proposed framework does not eliminate risk, particularly in environments with unmodeled hazards or adversarial conditions, and should be complemented by human oversight and additional safety mechanisms.
+We release our benchmarks and evaluation protocols to support transparent and reproducible research on safe adaptation under distribution shift.
 Appendix A Algorithm
-1:Input: discount factor $\gamma$, cost threshold $\delta$, number of training steps $T_{\max}$, batch size NN, env time limit $t_{\max}$, episodes-per-history range $K_{\min},K_{\max}]$, CTG range $\mathrm{CTG}_{\min},\mathrm{CTG}_{\max}]$, environment distribution $\mathcal{E}$
+Algorithm 1 SCARED Implemented with DDPG
+1:Input: discount factor $\gamma$, cost budget $\delta$, number of training steps $T_{\max}$, batch size NN, env time limit $t_{\max}$, episodes-per-history range $K_{\min},K_{\max}]$, CTG range $\mathrm{CTG}_{\min},\mathrm{CTG}_{\max}]$, environment distribution $\mathcal{E}$
 2:Initialize: actor $\pi(\cdot\mid s,H,\mathrm{CTG};\theta_{p})$, reward critic $Q(s,a;\theta_{v})$, cost critic $Q^{c}(s,a;\theta_{c})$, and target nets $\{\pi^{\prime},Q^{\prime},Q^{c\prime}\}$.
@@
 3:for $T\leftarrow 1$ to $T_{\max}$ do
-4:  $K\leftarrow\texttt{rnd}(K_{\min},K_{\max})$,  $\mathrm{CTG}\leftarrow\texttt{rnd}(\mathrm{CTG}_{\min},\mathrm{CTG}_{\max})$,  $t\leftarrow 0$
-$H\leftarrow[\;]$ // list of episodes
+4:  $K\leftarrow\texttt{rnd}(K_{\min},K_{\max})$,  $\mathrm{CTG}\leftarrow\texttt{rnd}(\mathrm{CTG}_{\min},\mathrm{CTG}_{\max})$,  $t\leftarrow 0$ $H\leftarrow[\;]$ // list of episodes
 […]
@@
 34:   end for
-35:   $d\lambda\leftarrow d\lambda+C_{e}^{\max}$
+35:   $d\lambda\leftarrow d\lambda+(C_{e}^{\max}-\delta)$
 36:  end for
@@
 39:end for
-Algorithm 1 EPPO Implemented with DDPG
 Note that our proposed optimization (6) can be used with any policy optimizer.
@@
 Proof.
-Let $J(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}$ in this proof.
+Let $J(\pi)\coloneqq{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}$ in this proof.
 We prove the theorem in both directions.
@@
 Subtract this from $J(\pi)$ and combine with (7):
-$L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big)\;=\;J(\pi)-\|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi)\;\leq\;J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\;\leq\;p^{\star}.$ (9)
+$L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big)\;$ $=\;J(\pi)-\|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi)$
+$\leq\;J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)$
+$\leq\;p^{\star}.$ (8)
 For a primal-optimal $\pi^{\star}$, feasibility gives $\sum_{i}g_{i}^{+}(\pi^{\star})=0$, hence
 $L_{\Sigma}\big(\pi^{\star},\ \|\lambda^{\star}\|_{\infty}\big)=J(\pi^{\star})=p^{\star}.$
-Together with (9), this shows
+Together with (B.1), this shows
 $\pi^{\star}\in\arg\max_{\pi}L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big).$
@@
 If constraints are strict ($s(\pi_{t+1})<0$), then $\lambda_{t+1}\leq\lambda_{t}$.
-Under persistent violation, $\lambda_{t}$ eventually exceeds $\|\lambda^{*}\|_{\infty}$, by the exact-penalty bound (9), any maximizer of $L_{\Sigma}(\cdot,\lambda_{t})$ is then feasible/optimal and $s(\pi_{t+1})\leq 0$, so $\lambda$ stabilizes.
+Under persistent violation, $\lambda_{t}$ eventually exceeds $\|\lambda^{*}\|_{\infty}$, by the exact-penalty bound (B.1), any maximizer of $L_{\Sigma}(\cdot,\lambda_{t})$ is then feasible/optimal and $s(\pi_{t+1})\leq 0$, so $\lambda$ stabilizes.
 ∎
+Note: The full convergence proof to an exact fixed point is beyond the scope of this work.
 B.2 Proof of Proposition 1
@@
 Then, $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$ and $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$ implies:
-$\mathbb{P}_{\text{train}}((i,j))=\frac{e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}}\quad\text{and}\quad\mathbb{P}_{\text{test}}((i,j))=\frac{e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{\alpha}},$
+[…]
+$\mathbb{P}_{\text{test}}((i,j))=\frac{e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{\alpha}}$
 […]
@@
 We keep the same notation to the proof so far. By the definition of KL divergence and explicit form of the probabilities, we have
-$D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)$ $=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{\mathbb{P}_{\text{train }}((i,j))}{\mathbb{P}_{\text{test }}((i,j))}$ (12)
-$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{Z_{\alpha}e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}$ (13)
+$D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)$
+$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{\mathbb{P}_{\text{train }}((i,j))}{\mathbb{P}_{\text{test }}((i,j))}$
+$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{Z_{\alpha}e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}$
 $=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\left(\log\frac{Z_{\alpha}}{Z_{-\alpha}}-2\alpha((i-i_{c})^{2}+(j-j_{c})^{2})\right).$ (11)
@@
 In our experiments, we set $n=5$.
-Reinforcement Pretraining.
+For SafeVelocity, we use the Ant and HalfCheetah environments from SafetyGym (Ji et al., 2023), but sample a different target velocity for each trajectory to convert them into safe ICRL benchmarks.
+Other aspects of the environments remain unchanged.
+SCARED.
 While running Algorithm 1 for reinforcement pretraining, we resample a new environment from the training distribution described in Section 5 every $K$ episodes.
@@
 A prediction head outputs either the action distribution (for discrete actions) or the value (for continuous actions).
-In practice, since the environments are generated randomly, it is possible that an environment doesn’t have a feasible solution, causing the agent to inevitably violate the cost threshold.
-This results in $\lambda$ growing too large in an attempt to mitigate these cases.
-To avoid this, we suggest capping $\lambda$ to a reasonably large value.
-Ignoring abnormally high-cost episodes is another solution worth exploring.
-Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
-$K_{\min}$, $K_{\max}$ 50, 50 20, 20 20, 20
-Episode time limit $t_{\max}$ 30 75 75
-Replay buffer capacity 100,000 100,000 100,000
-Embedding Dim 64 64 64
-Hidden Dim 64 64 64
-Num Layers 4 4 4
-Num Heads 8 8 8
-Seq Len 1500 1500 1500
-Attention Dropout 0 0 0
-Residual Dropout 0 0 0
-Embedding Dropout 5 5 5
-Learning Rate 3e-4 3e-4 3e-4
-Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
-Clip Grad 1.0 1.0 1.0
-Batch Size 32 32 32
-Num Updates 30k 10k 10k
-Optimizer Adam Adam Adam
-Table 1: Parameters for Safe Reinforcement Pretraining
-Supervised Pretraining.
-In supervised pretraining, we collect a dataset $\mathcal{D}=\quantity{\Xi_{i}}$ comprising multiple trajectories, where each trajectory $\Xi_{i}\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$ represents a sequence of episodes generated by running existing safe RL algorithms on various CMDPs.
+Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
+$K_{\min}$, $K_{\max}$ 50, 50 20, 20
+Episode time limit $t_{\max}$ 30 75
+Replay buffer capacity 100,000 100,000
+Embedding Dim 64 64
+Hidden Dim 64 64
+Num Layers 4 4
+Num Heads 8 8
+Seq Len 1500 1500
+Attention Dropout 0 0
+Residual Dropout 0 0
+Embedding Dropout 5 5
+Learning Rate 3e-4 3e-4
+Betas (0.9, 0.99) (0.9, 0.99)
+Clip Grad 1.0 1.0
+Batch Size 32 32
+Num Updates 30k 10k
+Optimizer Adam Adam
+Table 1: Parameters for SCARED
+SafeMeta & MAML with penalty
+We use the original implementations by (Xu & Zhu, 2025) with 3 hidden layers of (64, 512, 64) units for policy, value, and cost networks, totaling 205,510 parameters, comparable to SCARED.
+We train for 15k meta-iterations, where each iteration samples 20 tasks from the training distribution.
+During meta-testing, both algorithms perform $K$ gradient updates to adapt to new tasks with unseen cost-to-go values.
+Note that while some meta-RL methods use frame stacking or recurrent architectures, their hidden states reset between episodes, making them fundamentally episodic—they do not carry information across episode boundaries as ICRL methods do.
+Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
+$K_{\min}$, $K_{\max}$ 50, 50 20, 20
+Episode time limit $t_{\max}$ 30 75
+Hidden layers (64, 512, 64) (64, 512, 64)
+Min/Max batch size 500/1500 1500/1500
+Policy learning rate 1e-3 1e-3
+Value/Cost learning rate 3e-2/1e-1 3e-2/1e-1
+Discount factor $\gamma$ 0.99 0.99
+GAE parameter $\tau$ 0.95 0.95
+Max KL divergence 1e-3 1e-3
+Lagrangian weight $\lambda$ 1.0 1.0
+Num meta-training iterations 15k 15k
+Policy optimizer Adam Adam
+Value/Cost optimizer L-BFGS L-BFGS
+Table 2: Parameters for SafeMeta and MAML with penalty
+Safe Algorithm Distillation.
+In safe algorithm distillation, we collect a dataset $\mathcal{D}=\quantity{\Xi_{i}}$ comprising multiple trajectories, where each trajectory $\Xi_{i}\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$ represents a sequence of episodes generated by running existing safe RL algorithms on various CMDPs.
 Each episode $\tau_{k}$ in a trajectory $\Xi_{i}$ comprises states, actions, rewards, and costs, with the episode return $G(\tau_{k})\doteq\sum_{t=1}^{T}R_{t}$ and cost-to-go $G_{c,t}(\tau_{k})\doteq\sum_{i=t+1}^{T}C_{i}$, expected to increase and decrease, respectively, with $k$ as the RL algorithm learns.
-The policy $\pi_{\theta}$ is trained autoregressively using an imitation learning loss to distill the behavior of RL algorithms demonstrated in the dataset, following the concept of algorithm distillation (Laskin et al., 2023).
-For discrete action spaces (e.g., SafeDarkRoom), we use a cross-entropy loss, defined as:
-$\mathcal{L}_{\mathrm{CE}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[-\log\pi_{\theta}(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right],$ (16)
-where $\pi_{\theta}$ is the categorical distribution over discrete actions.
-For continuous action spaces (e.g., SafeDarkMujoco), we use an L2 loss, where the transformer outputs an action mean $\mu_{t}=\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))$, minimizing:
-$\mathcal{L}_{\mathrm{L2}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[\left\|A_{t}^{k}-\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right\|^{2}\right].$ (17)
-These losses extend the imitation learning loss from Equation (5) by incorporating RTG and CTG conditioning, aligning with our goal of distilling RL algorithms into the policy’s forward pass while optimizing for safety constraints.
+We train the policy $\pi_{\theta}$ autoregressively using safe algorithm distillation to distill the behavior of safe RL algorithms present in the dataset, following Laskin et al.
+(2023).
+Concretely, each training example consists of a trajectory segment $\Xi_{i}$ containing state and action pairs along with RTG and CTG conditioning.
+Let the policy take as input $(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))$ and predict the action at time $t$ for episode $k$.
+Discrete action spaces.
+For environments with discrete actions (e.g., SafeDarkRoom), the model outputs categorical logits $z_{\theta}(\cdot)$, generating a categorical distribution $\pi_{\theta}(\cdot\mid\cdot)=\mathrm{Softmax}(z_{\theta}(\cdot))$.
+We train with the standard cross-entropy loss:
+$\mathcal{L}_{\mathrm{disc}}(\theta)$ $=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[-\log\pi_{\theta}\!\left(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)\right].$ (13)
+Continuous action spaces.
+For environments with continuous actions (e.g., SafeDarkMujoco), the model outputs a conditional distribution over actions parameterized by a mean $\mu_{\theta}(\cdot)$.
+We supervise the predicted mean using an $\ell_{2}$ loss:
+$\mathcal{L}_{\mathrm{cont}}(\theta)$ $=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[\left\|A_{t}^{k}-\mu_{\theta}\!\left(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)\right\|_{2}^{2}\right].$ (14)
+These objectives enables the transformer to distill constraint-aware goal seeking behaviors into its forward pass.
 Dataset Collection.
-For supervised pretraining, we collect learning trajectories using reinforcement learning (RL) algorithms.
-As our base RL algorithm, we employ PPO-Lagrangian Schulman et al.
-(2017); Ray et al.
-(2019), which is designed to maximize rewards while enforcing safety constraints.
+For safe AD, we collect learning trajectories using safe RL algorithms.
+As our base safe RL algorithm, we employ PPO-Lagrangian (Schulman et al., 2017; Ray et al., 2019), which is designed to maximize rewards while enforcing safety constraints.
 These trajectories capture behaviors that learn to avoid obstacles.
@@
 For the SafeDarkRoom environment, we use three cost limits: 0, 2.5, and 5.0.
-For the SafeDarkMujoco environment, we similarly use three cost limits: 0, 5, and 10.
-For each cost limit, we collect 50,000 steps of learning history in both environments.
+For each cost limit, we collect 50,000 steps of learning history.
+For the SafeDarkMujoco and SafetyVelocity environments, prior offline RL benchmarks and constrained decision transformer methods typically use datasets on the order of 1122 million environment steps (Fu et al., 2020; Liu et al., 2023).
+In our experiments, we intentionally use larger datasets of 2255 million steps.
+This difference from the usual convention is motivated by the need to collect diverse cost-related experiences in continuous control settings, where meaningful safety-constrained behavior emerges only after long training horizons.
+In our data collection, we set a single cost limit of 0 throughout the full learning history.
+This setup intentionally includes trajectories generating a various range of cost violations, from high-cost exploration phases to near-zero-cost behavior as training progresses.
+Continuous control tasks are substantially more difficult to solve than discrete ones, and learning meaningful policies under safety constraints typically requires more interaction data.
+Thus, this design poses a challenge for safe supervised in-context RL: conditioning on both return-to-go and cost-to-go induces a large combinatorial space of reward–cost trade-offs that must be represented in the dataset.
+As a result, the required dataset size grows substantially.
 Hyper Parameters.
-We report the hyperparameters used for reinforcement pretraining (Table 1) and supervised pretraining (Table 2).
-Reinforcement pretraining adopts the AMAGO framework (Grigsby et al., 2024a) integrated with our novel EPPO method.
-Supervised pretraining employs a constrained decision transformer as the backbone (Liu et al., 2023).
-Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
-Embedding Dim 64 64 64
-Hidden Dim 512 256 256
-Num Layers 8 8 8
-Num Heads 8 8 8
-Seq Len 100 300 200
-Attention Dropout 0.5 0.5 0.5
-Residual Dropout 0.1 0.1 0.1
-Embedding Dropout 0.3 0.3 0.3
-Learning Rate 3e-4 3e-4 3e-4
-Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
-Clip Grad 1.0 1.0 1.0
-Batch Size 512 128 128
-Num Updates 300k 500k 500k
-Optimizer Adam Adam Adam
-Table 2: Parameters for Safe Supervised Pretraining
-Appendix D Variation of Supervised Pretraining
-Figure 4: Performance Comparison of Algorithm Distillation (AD) style supervised pretraining and Algorithm Distillation with Noise (AD-EPS) style supervised pretraining in SafeDarkRoom.
-AD-EPS fails to generalize in OOD environments.
-In this section, we compare the performance of Algorithm Distillation (AD) (Laskin et al., 2023) and Algorithm Distillation with Noise (AD-EPS) (Zisman et al., 2023) in SafeDarkRoom.
-AD-EPS is designed to learn in-context RL algorithms using datasets generated from a perturbed optimal policy.
-This approach allows efficient generation of learning trajectories from a single optimal policy.
-However, our results in complex environments with cost signals reveal that AD-EPS relies on artificial trajectories.
-AD-EPS fails to learn effective in-context RL algorithms when relying solely on perturbed optimal policies, as the perturbations do not account for obstacle avoidance in SafeDarkRoom environment and instead introduce random action variations.
+We report the hyperparameters used for SCARED (Table 1) and safe AD (Table 3).
+SCARED adopts the AMAGO framework (Grigsby et al., 2024a) integrated with our SCARED method.
+Safe AD employs a constrained decision transformer as the backbone (Liu et al., 2023).
+Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
+Embedding Dim 64 512
+Hidden Dim 512 256
+Num Layers 8 8
+Num Heads 8 8
+Seq Len 100 200
+Attention Dropout 0.5 0.5
+Residual Dropout 0.1 0.1
+Embedding Dropout 0.3 0.3
+Learning Rate 3e-4 3e-4
+Betas (0.9, 0.99) (0.9, 0.99)
+Clip Grad 1.0 1.0
+Batch Size 512 128
+Num Updates 300k 500k
+Optimizer Adam Adam
+Table 3: Parameters for Safe Algorithm Distillation
+Appendix D Variation of Safe Algorithm Distillation
+Figure 4: Performance comparison between algorithm distillation and algorithm distillation with noise in SafeDarkRoom. AD-EPS fails to generalize to out-of-distribution (OOD) environments.
+In this section, we compare safe algorithm distillation (AD) (Laskin et al., 2023) and safe algorithm distillation with noise (AD-EPS) (Zisman et al., 2023) in SafeDarkRoom environment.
+AD-EPS aims to learn in-context RL algorithms by training on datasets generated from a single optimal policy with injected action noise, enabling efficient collection of learning trajectories.
+However, in environments with safety constraints and explicit cost signals, we find that AD-EPS struggles to generalize.
+In SafeDarkRoom, the perturbed trajectories produced by AD-EPS fail to capture meaningful obstacle-avoidance behavior, as the injected noise primarily induces random action variations rather than structured exploration.
+As a result, AD-EPS relies on artificial trajectories that do not reflect the true learning dynamics required for effective in-context RL, leading to poor generalization in out-of-distribution settings.
 Appendix E Ablation Studies
-In this section, we present our ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, as well as specific factors like dataset size for supervised pretraining.
+In this section, we present our ablation studies on SCARED and safe AD, examining shared factors such as context length and model size, as well as specific factors like dataset size for safe AD.
 Our findings reveal distinct sensitivities to these factors.
-Reinforcement pretraining remains largely unaffected by model size, whereas supervised pretraining performance is significantly influenced by model size (See Figures 5.(b) and 6.(b)).
-Regarding context length, reinforcement pretraining benefits from longer sequences, while supervised pretraining exhibits degraded performance with longer contexts.
-Conversely, supervised pretraining performs better with shorter context lengths, where reinforcement pretraining struggles (See Figures 5.(a) and 6.(a)).
+SCARED remains largely unaffected by model size, whereas safe AD performance is significantly influenced by model size (See Figures 5.(b) and 6.(b)).
+Regarding context length, SCARED benefits from longer sequences, while safe AD shows degraded performance with longer contexts.
+Conversely, safe AD performs better with shorter context lengths, where SCARED struggles (See Figures 5.(a) and 6.(a)).
 These results indicate that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, manages long-term credit assignment more effectively.
-We also confirm the well-established finding that supervised pretraining is highly sensitive to dataset size, with models failing to learn and exhibiting random behavior on out-of-distribution data when the dataset is small (See Figure 6.(c)).
+We also confirm the well-established finding that safe AD is highly sensitive to dataset size, with models failing to learn and exhibiting random behavior on out-of-distribution data when the dataset is small (See Figure 6.(c)).
 (a) Context Length
@@
 (c) Optimization
-Figure 5: Ablations of Safe Reinforcement Pretraining on SafeDarkRoom.
+Figure 5: Ablations of SCARED on SafeDarkRoom.
 (a), (b) The evaluation is set up similarly to Question (i).
@@
 (c) At each training step, the average total return of 50 episodes across 10 random test environments, and the average Q Value across 50 episodes of 250 random train enironments are plotted.
-EPPO is easier to tune and more stable to train than the naive primal-optimal method.
+Our SCARED optimization is easier to tune and more stable to train than the naive primal-optimal method.
 (a) Context Length
@@
 (c) Dataset Size
-Figure 6: Ablations of Safe Supervised Pretraining on SafeDarkRoom.
+Figure 6: Ablations of Safe Algorithm Distillation on SafeDarkRoom.
 The evaluation is set up similarly to Question (i).